### v2

# package/hyperref/before

Abstract
In this paper, we propose a neural multiscale decomposition method (NeuralMD) for solving the nonlinear Klein-Gordon equation (NKGE) with a dimensionless parameter $\varepsilon\in(0,1]$ from the relativistic regime to the nonrelativistic limit regime. The solution of the NKGE propagates waves with wavelength at $O(1)$ and $O(\varepsilon^{2})$ in space and time, respectively, which brings the oscillation in time. Existing collocation-based methods for solving this equation lead to spectral bias and propagation failure. To mitigate the spectral bias induced by high-frequency time oscillation, we employ a multiscale time integrator (MTI) to absorb the time oscillation into the phase. This decomposes the NKGE into a nonlinear Schrödinger equation with wave operator (NLSW) with well-prepared initial data and a remainder equation with small initial data. As $\varepsilon\to 0$, the NKGE converges to the NLSW at rate $O(\varepsilon^{2})$, and the contribution of the remainder equation becomes negligible. Furthermore, to alleviate propagation failure caused by medium-frequency time oscillation, we propose a gated gradient correlation correction strategy to enforce temporal coherence in collocation-based methods. As a result, the approximation of the remainder term is no longer affected by propagation failure. Comparative experiments with existing collocation-based methods demonstrate the superior performance of our method for solving the NKGE with various regularities of initial data over the whole regime.
keywords:
Neural multiscale decomposition, Nonlinear Klein-Gordon equation, Nonrelativistic limit regime, Spectral bias, Propagation failure.
{AMS}
65N22, 65N55, 68T07
1 Introduction
The dimensionless nonlinear Klein–Gordon equation (NKGE) with cubic nonlinearity in $d$-dimensions ($d=1,2,3$) [bao2014mti, bao2012analy, grundland1992, messiah2014] is given by
(1) $\left\{\begin{aligned} &\varepsilon^{2}\partial_{tt}u(\mathbf{x},t)-\Delta u(\mathbf{x},t)+\varepsilon^{-2}u(\mathbf{x},t)+\lambda|u(\mathbf{x},t)|^{2}u(\mathbf{x},t)=0,\mathbf{x}\in\mathbb{R}^{d},\ t>0,\\[5.69054pt] &u(\mathbf{x},0)=\phi_{1}(\mathbf{x}),\qquad\partial_{t}u(\mathbf{x},0)=\varepsilon^{-2}\phi_{2}(\mathbf{x}),\qquad\mathbf{x}\in\mathbb{R}^{d},\end{aligned}\right$.
where $u=u(\mathbf{x},t)$ is a complex-valued field, $0<\varepsilon\leq 1$ is a dimensionless parameter inversely proportional to the speed of light, and $\lambda\in\mathbb{R}$ characterizes the nonlinear interaction (defocusing for $\lambda>0$ and focusing for $\lambda<0$). $\phi_{1}$ and $\phi_{2}$ denote prescribed $\varepsilon$-independent initial data.
When $\varepsilon=1$ in (1), corresponding to the $O(1)$-wave-speed regime, various numerical methods for the associated Cauchy problem have been proposed and analyzed [bratsos2009numer, duncan1997symp, liu2018sym]. In particular, finite difference time domain (FDTD) methods [duncan1997symp, jimenez1990, strauss1978] are highly efficient and accurate in this regime. In contrast, in the nonrelativistic limit $0<\varepsilon\ll 1$, the design and analysis of efficient and accurate numerical methods become significantly more challenging due to the highly oscillatory behavior of the solution in time. To address this issue, a variety of uniformly accurate (UA) schemes have been developed for the nonrelativistic limit regime of (1), whose error bounds are uniform with respect to $\varepsilon\in(0,1]$ and thus permit time steps independent of the fast temporal scale $O(\varepsilon^{2})$ [bao2012analy]. These include the multiscale time integrator (MTI) [bao2014mti, bao2017uniform, bao2019comp], the two-scale formulation (TSF) method [chartier2015], and the nested Picard iterative integrator (NPI) [cai2021uni, cai2022uni, li2025uniform], all of which achieve super-resolution in time for highly oscillatory solutions in the nonrelativistic limit.
However, these schemes still provide the solution in a fashion that is local in time and rely on refined temporal discretizations, particularly in the nonrelativistic limit regime. For simulations over long time intervals, this leads to a substantial computational burden. Moreover, for each new final time or spatial domain, the full time-dependent solution has to be recomputed on a new grid in time and space. To address this issue, a class of physics-driven collocation-based methods [raissi2019phy, wang2023sci, cuomo2022sci] has been proposed, which does not rely on precomputed numerical data, avoids time-step error accumulation, and aims to obtain the solution over the entire time interval in a single computation. However, when applied to the NKGE with pronounced time oscillation, existing methods typically suffer from spectral bias [rahaman2019] and propagation failure [wang2022resp], and they may completely break down as $\varepsilon\to 0$.
For spectral bias, recent studies have shown that deep neural networks trained by gradient-based methods tend to learn low-frequency components of the target function much faster than high-frequency components, a phenomenon often referred to as the frequency principle [xu2019freq, rahaman2019, xu2025view]. This preferential learning is particularly problematic when the underlying PDE solution exhibits a wide range of active scales, as high-frequency modes are learned only at later stages of training or may not be captured at all [luo2022upper]. Consequently, neural-network-based solvers may fail to resolve sharp layers, oscillatory structures, and multiscale patterns, even when the network has sufficient approximation capacity in principle [xu2019freq, rahaman2019]. This issue is further exacerbated in high-dimensional, stochastic, or geometrically complex problems, where spectral bias interacts with sampling and optimization challenges, leading to significant degradation of solution quality [xu2025view]. These findings highlight the need for architectures and training strategies that explicitly mitigate spectral bias to improve the performance of neural PDE solvers for multiscale problems [xu2019freq, luo2022upper]. Recent works have identified that PINNs struggle to converge to solutions when target functions exhibit high-frequency patterns [krishnapriyan2021, moseley2023finite]. The application of differential operators to neural networks complicates the loss landscape and makes optimization more difficult. Analyses based on the neural tangent kernel (NTK) [wang2022why, farhani2022mom] show that components associated with larger eigenvalues converge more slowly, reinforcing spectral bias, while existing remedies such as adaptive loss reweighting and momentum-based optimizers like Adam provide only limited improvements for highly oscillatory targets. An effective strategy involves applying Fourier feature mapping to the input coordinates, which has been shown to successfully capture high-frequency patterns in various PDEs [hertz2021sape, tancik2020fourier, li2023deep].
For propagation failure, existing collocation-based methods have shown systematic breakdowns in training, even for relatively simple PDEs, leading to qualitatively incorrect solutions despite apparently successful optimization [krishnapri2021]. One line of work attributes the failure modes to optimization difficulties with new sampling strategies [wu2023comp, wang2023r3], imbalances among loss terms [yu2022grad, wu2024ropinn, wu2025propinn], optimizers [rathore2024], automatic differential methods [shi2024stoch], and sensitivity to PDE coefficients [wang2021under, wang2022why] etc. Enforcing only the equation residual in the interior region can cause propagation failure, where small residual loss and large approximation error arise due to insufficient supervision from initial and boundary data to interior collocation points [wang2023r3]. To alleviate this effect, several approaches refine the distribution of collocation points by adaptively concentrating them in regions with large residuals, or by iteratively augmenting the training set based on loss indicators [lu2021deepxde, nabian2021]. Other works have proposed modified loss formulations, higher-order norms for the PDE residual, and alternative optimization procedures, but these remedies mainly target training stability rather than the structural origins of propagation failure [yu2022grad, wu2024ropinn, wang2022l2, rathore2024, shi2024stoch]. In contrast, existing methods treat collocation points as independent samples, limiting the propagation of boundary data and making the network prone to trivial solutions and local minima [wong2022learn, rohrhofer2022, rohrhofer2022role]. In addition, causality-aware training procedures for time-dependent PDEs have shown that explicitly encoding directional information flow in time can mitigate propagation failures and improve the reliability [wang2022resp].
Motivated by the aforementioned challenges and existing methods, we propose a neural multiscale decomposition method, termed NeuralMD, to address the spectral bias and propagation failure induced by time oscillation in the NKGE. Our main contributions can be summarized as follows:
  • We design NeuralMD by decomposing the solution of NKGE into explicit high-frequency phase factors and a low-frequency envelope governed by an amplitude equation. This decomposition is based on a global-in-time WKB expansion with remainder terms, which splits the NKGE into a nonlinear Schrödinger equation with a wave operator (NLSW) with well-prepared initial data, and a remainder equation with small initial data. NeuralMD drops the high-frequency time oscillation into the phase, requiring only solving the lightly oscillating NLSW in the nonrelativistic limit regime. Compared to existing methods that mitigate spectral bias, NeuralMD significantly reduces the difficulty of solving NKGE.
  • NeuralMD is a two-stage pretraining framework. The first stage trains a network to approximate the modulated NLSW, and the second stage trains a remainder network to correct amplitude errors. As $\varepsilon\to 0$, we solve only the NLSW and recover the oscillatory NKGE solution via a WKB expansion without remainder terms. As $\varepsilon\to 1$, we also solve the remainder equation and use a WKB expansion with remainder terms. For intermediate $\varepsilon$, a WKB reconstruction–based criterion decides whether the remainder is included, so that NeuralMD applies uniformly across the whole regime without manual multiscale partitioning.
  • Neglecting the light time oscillation of the modulated NLSW has little effect on the reconstruction, whereas the remainder equation inherits the strong time oscillation of the NKGE. To alleviate premature propagation failure when solving the remainder, we propose a gated gradient correlation correction strategy to enforce temporal coherence among nearby collocation points. Concretely, we introduce random temporal perturbations to form local neighborhoods and correlate residuals across perturbed time steps, dynamically removing collocation points in low-residual regions and reallocating samples to high-residual regions.
  • Furthermore, we extend NeuralMD into an interpretable version by introducing Kolmogorov–Arnold networks (KANs) [liu2024kan], which provide structural interpretability for oscillation dropping and remainder amplitude compensation. We evaluate the effectiveness and robustness of NeuralMD on the NKGE with different initial data regularities, testing both problems over the global time domain and those extending beyond long time intervals. Numerical experiments show that NeuralMD effectively mitigates spectral bias and propagation failure induced by time oscillation, achieving high accuracy across the full range of $\varepsilon\in(0,1]$.
The remainder of the paper is organized as follows. Section 2 introduces the phenomenon of spectral bias and propagation failure induced by the time oscillation of the NKGE. Section 3 presents the NeuralMD method and discusses its implementation. Section 4 showcases numerical results. Finally, Section 5 introduces the conclusion and outlook.
2 Main Problem
2.1 Oscillation in time induces spectral bias
For (1), as $\varepsilon\to 0$, the solution develops rapidly oscillatory phases in time with carrier frequencies $\pm 1/\varepsilon^{2}$, so that the temporal wavelength is only $O(\varepsilon^{2})$ while the spatial scale remains $O(1)$. This makes the problem extremely stiff in time. In the same regime, the conserved energy satisfies $E(t)=O(\varepsilon^{-2})$ and becomes unbounded as $\varepsilon\to 0$. This “energy inflation” significantly complicates both asymptotic analysis and error control, since any discretization error is amplified at this energy scale and interacts with the fast phase, rendering long-time integration highly fragile.
We now investigate the challenge of spectral bias induced by time oscillation when solving this equation using collocation-based optimization methods such as PINNs. The model $u_{\theta}$ approximates the NKGE solution $u$ by minimizing the following loss function, including the residual, the initial condition, and the period boundary losses
(2) $\left\{\begin{aligned} \mathcal{L}_{\mathrm{Res}}(\theta)&=\mathbb{E}_{(\mathbf{x},t)\in P_{f}}\left\|\varepsilon^{2}\partial_{tt}u_{\theta}-\Delta u_{\theta}+\varepsilon^{-2}u_{\theta}+\lambda|u_{\theta}|^{2}u_{\theta}\right\|_{L^{2}}^{2},\\[4.0pt] \mathcal{L}_{\mathrm{Ic}}(\theta)&=\mathbb{E}_{(\mathbf{x},t)\in P_{0}}\left(\left\|u_{\theta}(x_{0},0)-\phi_{1}\right\|_{L^{2}}^{2}+\left\|\partial_{t}u_{\theta}(x_{0},0)-\varepsilon^{-2}\phi_{2}\right\|_{L^{2}}^{2}\right),\\[4.0pt] \mathcal{L}_{\mathrm{Bd}}(\theta)&=\mathbb{E}_{(\mathbf{x},t)\in P_{b}}\left(\left\|u_{\theta}(a,t_{a})-u_{\theta}(b,t_{b})\right\|_{L^{2}}^{2}+\left\|\partial_{x}u_{\theta}(a,t_{a})-\partial_{x}u_{\theta}(b,t_{b})\right\|_{L^{2}}^{2}\right).\end{aligned}\right$.
where $P_{f}$, $P_{0}$, and $P_{b}$ denote the sets of collocation points for the PDE residual, the initial condition, and the boundary condition, respectively. $\mathbb{E}_{(\mathbf{x},t)\in P}$ denotes the empirical average over the set $P$. $[a,b]$ is the period spatial domain.
Here, we introduce the following composition [dong2021period]
(3) $\begin{bmatrix}t&1&\cos\left(\frac{2\pi}{P}x\right)&\sin\left(\frac{2\pi}{P}x\right)&\cos\left(\frac{2\pi}{P}2x\right)&\cdots&\cos\left(\frac{2\pi}{P}mx\right)&\sin\left(\frac{2\pi}{P}mx\right)\end{bmatrix}^{\top}$,
where $m>0$ is a positive integer. A similar transformation can be applied to a higher-dimensional spatial domain $\mathbf{x}$. We then apply the following transformation to satisfy the initial condition automatically:
(4) $\tilde{u}_{\theta}=(1+t)e^{-t}u_{0}+te^{-t}\partial_{t}u_{0}+\left(1-e^{-t}-te^{-t}\right)u_{\theta},$
The total loss used for training via automatic differentiation and backpropagation is the following
(5) $\mathcal{L}(\theta)=\lambda_{\mathrm{Res}}\mathcal{L}_{\mathrm{Res}}(\theta)+\lambda_{\mathrm{Ic}}\mathcal{L}_{\mathrm{Ic}}(\theta)+\lambda_{\mathrm{Bd}}\mathcal{L}_{\mathrm{Bd}}(\theta).$
We analyze the impact of the time oscillation of NKGE on the training dynamics of PINNs. As an example, we consider the commonly used hyperbolic tangent activation function
(6) $\sigma(t)=\tanh(t)=\frac{e^{t}-e^{-t}}{e^{t}+e^{-t}},\quad t\in\mathbb{R}.$
and PINNs with one hidden layer, having $m$ neurons, a 1-dimensional input $t$, and a 1-dimensional output, i.e.,
(7) $h(t)=\sum_{j=1}^{m}a_{j}\sigma(w_{j}t+b_{j}),\quad a_{j},w_{j},b_{j}\in{\rm\mathbb{R}},$
where $\theta_{j}\triangleq\{w_{j},b_{j},a_{j}\}$, $w_{j}$, $a_{j}$, and $b_{j}$ are training parameters.
Then, the Fourier transform of $h(x)$ can be computed as follows
(8) $\hat{h}(k)=\sum_{j=1}^{m}\frac{2\pi a_{j}}{|w_{j}|}\exp\Big(\frac{b_{j}k}{w_{j}}\Big)\frac{1}{\exp(-\frac{\pi k}{2w_{j}})-\exp(\frac{\pi k}{2w_{j}})},$
where $k$ denotes the frequency.
By Parseval’s theorem [stein2011fourier], this loss is identical to the standard mean-squared error in the Fourier domain, that is
(9) $\mathcal{L}(\theta)=\int_{-\infty}^{+\infty}[\lambda_{\mathrm{Res}}\mathcal{L}_{\mathrm{Res}}(\theta,k)+\lambda_{\mathrm{Ic}}\mathcal{L}_{\mathrm{Ic}}(\theta,k)+\lambda_{\mathrm{Bd}}\mathcal{L}_{\mathrm{Bd}}(\theta,k))]d{k}.$
To study how gradient descent attenuates errors across frequencies, we examine the loss in the Fourier domain and its dependence on the frequency $k$. In the early stage of training, the weights typically satisfy $|w_{j}|\ll 1$, so the last term in (8) can be approximated by
(10) $\frac{1}{\exp(-\pi k/w_{j})-\exp(\pi k/w_{j})}\approx-\operatorname{sgn}(w_{j})\exp\left(-\pi\left|\frac{k}{w_{j}}\right|\right)=\exp\left(-\left|\frac{\pi k}{2w_{j}}\right|\right),$
Hence, the magnitude of the contribution from frequency $k$ to the gradient with respect to $\theta_{lj}$ is
(11) $\frac{\partial\mathcal{L}(\theta_{j},k)}{\partial\theta_{j}}\approx|\frac{\partial\mathcal{L}(\theta_{j},k)}{\partial w_{j}}|\exp\!\left(-\left|\frac{\pi k}{2w_{j}}\right|\right)F_{j}.$
[…]
(12) $\left|-\varepsilon^{2}k^{2}+\varepsilon^{-2}\right|,$
In the main oscillatory band $k\sim\varepsilon^{-2}$, we have $\varepsilon^{2}k^{2}\sim\varepsilon^{-2}$. Using this as the leading chain-rule factor, the dominant contribution to the gradient at temporal frequency $k$ with respect to $\theta_{lj}$ can be approximated by
(13) $\frac{\partial L(k)}{\partial\theta_{j}}\approx\bigl|-\varepsilon^{2}k^{2}+\varepsilon^{-2}\bigr|\,\bigl\lVert u_{\theta}(k)-u(k)\bigr\rVert\exp\!\left(-\frac{\pi|k|}{2|w_{j}|}\right)\,F_{j}.$
where $F_{lj}$ is an $O(1)$ factor incorporating the smoothing effects of the nonlinear term and the spatial derivatives on this frequency band.
As $\varepsilon\to 0$, the coefficient $\bigl|-\varepsilon^{2}k^{2}+\varepsilon^{-2}\bigr|$ grows like $O(\varepsilon^{-2})$, so the residual loss assigns very large nominal gradients to high temporal frequencies. At the same time, in the early stage of training, the network sensitivity at the same $k$ is exponentially suppressed by $\exp\!\left(-\frac{\pi|k|}{2|w_{j}|}\right)$. Consequently, for NKGE in the nonrelativistic limit regime, spectral bias is prone to break down. Fig. 1 shows that, in the pronounced nonrelativistic limit regime ($\varepsilon=0.1$), PINNs exhibit optimization stagnation and high-frequency drift in the presence of strong time oscillation.
Figure 1: The prediction solution of PINNs for $\varepsilon=0.1$.
2.2 Oscillation in time induces propagation failure
In this section, we further investigate whether a mild decrease in $\varepsilon$, and the resulting mild-frequency time oscillation already affect the propagation behavior of collocation-based optimization methods such as PINNs. The NKGE not only contains a potential term of size $\varepsilon^{-2}$ that induces high-frequency time oscillation, but, more importantly, the initial time derivative also scales like $\varepsilon^{-2}$. Thus, as $\varepsilon\to 0$, high-frequency difficulty is present from the very beginning of the evolution. When the initial gradient carries substantial high-frequency content, PINNs struggle to accurately propagate the corresponding high-frequency dynamics into the interior of the domain. The learned solution typically stalls after a short time and enters a propagation-failure regime. This failure is not merely a “frozen” solution, but rather stems from a mismatch between the global optimization objective used in PINNs training and the inherently local-in-time causal structure of time-dependent PDEs.
Fig. 1 shows a failure mode that is distinct from the spectral bias induced by spatial oscillation. Starting from well-prepared initial data, the solution evolves only up to about $t\approx 0.5$ and then ceases to propagate. This behavior is consistent with propagation failure. In this section, we explore how gradually increasing the oscillation frequency leads to propagation failure.
For the NKGE in the transition regime, the solution exhibits a temporal wavelength of $O(\varepsilon^{2})$, with mid-frequency time oscillation as $\varepsilon$ decreases. Both the potential term and initial time derivative scale like $\varepsilon^{-2}$, introducing substantial mid-frequency content from the start. In the PINN setting, the correct solution must propagate from initial/boundary points to interior collocation points. However, mid-frequency time oscillation complicates this propagation, making it difficult for the network to transmit the correct dynamics. As a result, some interior points may converge to trivial or low-frequency solutions before the correct solution can reach them. These trivial solutions then spread to nearby points, leading to large regions with incorrect solutions. Fig. 2 illustrates this propagation-failure mode of PINNs for a simple temporally oscillatory problem: as the number of training iterations increases, the optimization stagnates, and the solution ceases to propagate correctly. Thus, mid-frequency time oscillation in the transition regime, combined with the global optimization objective of PINNs, causes the training dynamics to fail in propagation.
Figure 2: PINNs for a simple temporally oscillatory problem.
For wave-type equations such as the NKGE, the time evolution is subject to causality and a finite propagation speed of information. The solution at a collocation point $(x_{i},t_{i})$ depends only on data in a certain region of the past. Traditional time-marching numerical schemes naturally respect this causal structure and effectively incorporate historical information. In contrast, PINNs perform optimization over the entire time domain at once, making it difficult to enforce the causal relationships among temporally distributed collocation points.
In the early stages of training, the network parameters $\theta$ are randomly initialized, but PINNs require the output $u_{\theta}$ to satisfy the initial condition at $t=0$ and the residual at $t=t_{i}$ (typically at the final time $T$ for global training). At this point, the neural network has not yet learned how to propagate the wave function’s dynamics from $t=0$ to $t=t_{i}$. Imposing the residual constraint too rigidly can cause premature optimization stagnation. This stagnation is not due to the failure of physical information propagation, but rather the failure of gradient information propagation during training. To quantify this, we define a stiffness coefficient based on the evolution of $u_{\theta}$ at times $t$ and $t+\delta t$
(14) $D_{\mathrm{PINNs}}\left(t,t+\delta t\right)=\lim_{\lambda\rightarrow 0}{\frac{\|u_{\theta}\left(\mathbf{x},t+\delta\right)-u_{\theta-\lambda\frac{\partial u_{\theta}}{\partial\theta}\big|_{t}}\left(\mathbf{x},t+\delta t\right)\|}{\lambda}},$
If $D_{\mathrm{PINNs}}\left(t,t+\delta t\right)<\epsilon$, this indicates that the effect of the time interval $(t,t+\delta t)$ on the model is less than an empirical threshold $\epsilon$, leading to stagnation of the gradient and triggering the propagation failure mode.
Further, we measure the stiffness coefficient of the above evolution failure mode using a gradient correlation metric, defined as
[…]
We consider that propagation failure occurs when the gradient correlation $G_{u_{\theta}}(t,t+\delta t)$ between two adjacent time points, $t$ and $t+\delta t$, is small. This indicates the failure of propagation between these two points in the time domain. We proceed to rewrite the (14) by employing a Taylor expansion centered at $(\theta,\mathbf{x}^{\prime})$,
(16) $\begin{split}D_{\mathrm{PINNs}}(t,t+\delta t)&=\lim_{\lambda\rightarrow 0}{\frac{\|u_{\theta}\left(\mathbf{x},t+\delta t\right)-u_{\theta-\lambda\frac{\partial u_{\theta}}{\partial\theta}\big|_{t}}(\mathbf{x},t+\delta t)\|}{\lambda}}\\ &=\lim_{\lambda\to 0}\frac{\bigg\|u(\mathbf{x},t+\delta t;\theta)-u({\mathbf{x},t+\delta t;\theta-\lambda\frac{\partial u_{\theta}}{\partial\theta}(\mathbf{x})})\bigg\|}{\lambda}\\ &=\lim_{\lambda\to 0}\frac{\bigg\|\bigg<\frac{\partial u_{\theta}}{\partial\theta}(\mathbf{x},t+\delta t),\lambda\frac{\partial u_{\theta}}{\partial\theta}(\mathbf{x},t)\bigg>+\mathcal{O}(\lambda^{2})\bigg\|}{\lambda}\\ &=\lim_{\lambda\to 0}\frac{\lambda G_{u_{\theta}}(t,t+\delta t)+\mathcal{O}(\lambda^{2})}{\lambda}\\ &=G_{u_{\theta}}(t,t+\delta t).\end{split}$
Since the functions $D_{\mathrm{PINNs}}(t,t+\delta t)$ and $G_{u_{\theta}}(t,t+\delta t)$ are equivalent, the smallness of $D$ ensures the smallness of $G$, and vice versa.
When $G_{u_{\theta}}\left(t,t+\delta t\right)=0$, the gradients between perturbation time steps $(t,t+\delta t)$ are orthogonal, indicating no correlation between them, which triggers the evolution failure mode. In the nonrelativistic limit regime as $\varepsilon\to 0$, at the same spatial location $x$, the gradient of the network parameters at perturbation time steps $(t,t+\delta t)$ often contains high-frequency components, so
(17) $G_{u_{\theta}}\left(\left(x,t\right),\left(x,t+\delta t\right)\right)\;\approx\;\int_{0}^{\infty}{S_{x}\left(k,\varepsilon\right)}\cos{\left(k\delta t\right)}dk,$
where $S_{x}\left(k,\varepsilon\right)\geq 0$ is the temporal angular frequency spectrum density of $\partial_{\theta}u_{\theta}$, given by
(18) $S_{x}\left(k,\varepsilon\right)\sim\frac{1}{\varepsilon}f\left(\frac{k}{\varepsilon}\right).$
In the nonrelativistic limit regime, high-frequency weights dominate, and for a given $\delta t$, they are more likely to cancel each other out, causing $G_{u_{\theta}}$ to decay faster and possibly even reverse sign. This means that the updated directions of the parameters at adjacent time steps become more orthogonal or cancel each other, making it harder to propagate correct information from the gradient at the initial time along the time axis to later times, thus triggering the evolution failure mode.
However, in the nonrelativistic limit regime, the time oscillation is severe, and the resulting evolution failure mode is often attributed to spectral bias. To exclude the influence of spectral bias, we select values of $\varepsilon$ in a mild time oscillation region, gradually increasing the final time and testing the impact of long-time behavior prediction on PINNs’ evolution failure. Although the time oscillation is mild, the distribution of $S_{x}(k)$ still depends on $\delta t$, and the impact of gradient initial values with time oscillation on the gradient direction accumulates over time. In long-time behavior prediction, this accumulated effect causes the gradient correlation between adjacent time steps to gradually decrease, eventually triggering the evolution failure mode. As $\varepsilon$ decreases, the accumulation effect accelerates, and the evolution failure mode is triggered more quickly. The gradient correlation decay satisfies
(19) $G_{u_{\theta}}\left(x,\delta t\right)\approx A\left(x\right)-\frac{1}{2}A\left(x\right)\left\langle k^{2}\right\rangle_{S,x}\delta t^{2}+O\left(\delta t^{4}\right),$
where $A\left(x\right)$ represents the norm of the parameter gradient, quantifying the gradient autocorrelation
(20) $A\left(x\right)=\int_{0}^{\infty}{S_{x}\left(k,\varepsilon\right)}dk=G_{u_{\theta}}(x,0)=\|\partial_{\theta}u_{\theta}\left(x,t\right)\|,$
$\left\langle k^{2}\right\rangle_{S,x}$ is the weighted second moment of high-frequency components, indicating the impact of high-frequency time oscillation on gradient decorrelation. It is expressed as
(21) $\left\langle k^{2}\right\rangle_{S,x}=\frac{\int_{0}^{\infty}{k^{2}S_{x}\left(k,\varepsilon\right)}dk}{\int_{0}^{\infty}{S_{x}\left(k,\varepsilon\right)}dk}\sim\frac{1}{\varepsilon^{2}}\int_{0}^{\infty}{k^{2}f\left(\frac{k}{\varepsilon}\right)}dk$
This leads to the following expression for $G_{u_{\theta}}$, as
(22) $G_{u_{\theta}}\left(x,\delta t\right)\approx A\left(x\right)-\frac{1}{2}A\left(x\right)\left(\frac{1}{\varepsilon^{2}}\int_{0}^{\infty}{k^{2}f\left(\frac{k}{\varepsilon}\right)}dk\right)\delta t^{2}+O\left(\delta t^{4}\right)$
As $\varepsilon$ decreases, $\left\langle k^{2}\right\rangle_{S,x}$ increases, causing the gradient correlation between adjacent time steps to decay faster. For small $\varepsilon$ near 1, the gradient correlation decays enough to trigger evolution failure, especially in long-time predictions, where this failure mode is more pronounced.
To assess the effect of time oscillation on the time evolution of collocation-based methods, we set $T=5$ and slightly reduce $\varepsilon$ to $0.6$, while keeping the network architecture and training hyperparameters fixed. The results in Fig. 3 show that, due to the random and isolated temporal collocation used by existing methods, temporal causality and recursive consistency are not explicitly enforced. As a result, phase errors introduced at different sampling times accumulate and gradually decorrelate, reducing the ability of the model to track the true solution. When $\varepsilon=0.6$, optimization stagnation appears near $t=2$, where both phase and amplitude errors grow and fail to recover to the correct trajectory.
Figure 3: The prediction solution of PINNs for $\varepsilon=0.6,t=5.0$.
For long-time prediction, with $T=10$ and the same value of $\varepsilon$, Fig. 4 shows that evolution failure appears near $t=3$ even in the non-oscillatory case $\varepsilon=1.0$. Time evolution curves at different spatial locations in Fig. 5 further illustrate that increasing the terminal time from $T=5$ to $T=10$ leads to progressive stagnation as $t$ advances. These results indicate that the lack of causal and recursive constraints between temporal collocation points is the essential cause of propagation failure, while time oscillation merely accelerates its onset.
Figure 4: The prediction solution of PINNs for $\varepsilon=1.0,T=10.0$.
(a) $\varepsilon=1.0,T=5.0$
(b) $\varepsilon=1.0,T=10.0$
Figure 5: The prediction solution of PINNs for $\varepsilon=1.0$ in different T.
3 Our proposed method
In this section, we first introduce the MTI method based on the WKB expansion, which motivates the multiscale decomposition idea used in our proposed method. This decomposition separates the high-frequency time oscillation into the phase, while the amplitude remains only mildly oscillatory. However, the WKB expansion focuses solely on the limiting behavior in the nonrelativistic regime, and MTI itself is a local-in-time uniformly accurate method. Although these techniques are well established, our contribution lies in introducing the MTI methodology into PINNs for the first time. Motivated by this, we develop the NeuralMD method to mitigate spectral bias and enable effective learning of high-frequency time oscillation, particularly in the nonrelativistic limit regime. We further introduce a gated gradient correlation strategy to enforce temporal coherence in PINNs and alleviate propagation failure induced by the mildly oscillatory modulation equation and the small-amplitude yet high-frequency remainder equation. Finally, we present an interpretable variant of NeuralMD together with the two-stage training framework used in the proposed method.
3.1 Multiscale decomposition by frequency
The WKB expansion (or termed modulated Fourier expansion, MFo) is a classical tool in the analysis and numerical treatment of oscillatory problems [cohen2003modul, faou2013sob, hairer2006]. It has recently been applied in [faou2014asym] as a numerical integrator for (1) in the nonrelativistic regime $\varepsilon\to 0$. The method represents the solution by separating the rapid time oscillation from the slowly varying amplitudes. The expansion takes the form
(23) $u(\mathbf{x},t)=\sum_{m\in\mathbb{Z}}e^{imt/\varepsilon^{2}}u_{m}(\mathbf{x},t),$
where the time derivatives of $u_{m}(\mathbf{x},t)$ remain uniformly bounded as $\varepsilon\to 0$ for sufficiently smooth solutions. Retaining only the leading-order mode $m=1$, the expansion reduces to (see [faou2014asym, masmoudi2002])
(24) $u(\mathbf{x},t)=e^{it/\varepsilon^{2}}z(\mathbf{x},t)+e^{-it/\varepsilon^{2}}\bar{z}(\mathbf{x},t)+o(\varepsilon^{2}),\quad\varepsilon\to 0.$
where $z(\mathbf{x},t)$ is complex-valued and $\overline{z}$ is its complex conjugate. Under well-prepared initial data, the NKGE formally reduces to the nonlinear Schrödinger equation with wave operator (NLSW) [bao2012uni, bao2014uni], is given by
(25) $\left\{\begin{split}&2i\partial_{t}z(\mathbf{x},t)+\varepsilon^{2}\partial_{tt}z(\mathbf{x},t)-\Delta z(\mathbf{x},t)+3\lambda|z(\mathbf{x},t)|^{2}z(\mathbf{x},t)=0,\mathbf{x}\in\mathbb{R}^{d},t>0,\\ &z(\mathbf{x},0)=\frac{1}{2}\left[\phi_{1}(\mathbf{x})-i\phi_{2}(\mathbf{x})\right]=:z_{0}(\mathbf{x}),\qquad\mathbf{x}\in\mathbb{R}^{d},\\ &\partial_{t}z(\mathbf{x},0)=\frac{i}{2}\left[-\Delta z_{0}(\mathbf{x})+3\lambda|z_{0}(\mathbf{x})|^{2}z_{0}(\mathbf{x})\right].\end{split}\right$.
By dropping the small term $\varepsilon^{2}\partial_{tt}z$ in (25), the model reduces to the limiting nonlinear Schrödinger equation (NLSE) [machihara2002, masmoudi2002]
(26) $\left\{\begin{split}&2i\partial_{t}z(\mathbf{x},t)-\Delta z(\mathbf{x},t)+3\lambda|z(\mathbf{x},t)|^{2}z(\mathbf{x},t)=0,\quad\mathbf{x}\in\mathbb{R}^{d},t>0,\\ &z(\mathbf{x},0)=\frac{1}{2}\left[\phi_{1}(\mathbf{x})-i\phi_{2}(\mathbf{x})\right]:=z_{0}(\mathbf{x}),\qquad\mathbf{x}\in\mathbb{R}^{d}.\end{split}\right$.
In the nonrelativistic regime, the asymptotic representations
(27) $\left\{\begin{split}u_{\rm nlsw}(x,t)&=e^{it/\varepsilon^{2}}z_{\rm nlsw}(x,t)+e^{-it/\varepsilon^{2}}\overline{z}_{\rm nlsw}(x,t),\\ u_{\rm nlse}(x,t)&=e^{it/\varepsilon^{2}}z_{\rm nlse}(x,t)+e^{-it/\varepsilon^{2}}\overline{z}_{\rm nlse}(x,t),\end{split}\right$.
motivate the corresponding error functions
(28) $\left\{\begin{split}\eta_{\rm nlsw}(t)&=\|u(\cdot,t)-u_{\rm nlsw}(\cdot,t)\|_{H^{1}},\\ \eta_{\rm nlse}(t)&=\|u(\cdot,t)-u_{\rm nlse}(\cdot,t)\|_{H^{1}}.\end{split}\right$.
The convergence of the NKGE with respect to $\varepsilon$ depends on the regularity of the initial data. For $\phi_{1},\phi_{2}\in H^{2}(\Omega)$, the solution converges uniformly in time to the NLSW. For $\phi_{1},\phi_{2}\in H^{3}(\Omega)$, it converges to the NLSE, though generally not uniformly in time. These estimates can be summarized as
(29a) $\|u(\cdot,t)-u_{\rm sw}(\cdot,t)\|_{H^{1}}$ $\leq C_{0}\,\varepsilon^{2},$ $t\geq 0,$ $\phi_{1},\phi_{2}\in H^{2}(\Omega),$
(29b) $\|u(\cdot,t)-u_{\rm s}(\cdot,t)\|_{H^{1}}$ $\leq(C_{1}+C_{2}T)\varepsilon^{2},$ $0\leq t\leq T,$ $\phi_{1},\phi_{2}\in H^{3}(\Omega).$
where the constants $C_{0}$, $C_{1}$, and $C_{2}$ are independent of $\varepsilon$.
Fig. 6 demonstrates that the time evolution of the error convergence curves agrees with the time uniform estimate in (29a). In addition, the results in Fig. 6 validate the time linear growth convergence behavior stated in (29b).
Figure 6: Error convergence curves under different $\varepsilon$. (a) The NKGE converges to the NLSW uniformly in time. (b) The NKGE converges to the NLSE with linear growth in time.
Inspired by the multiscale decomposition in (24), the MTI method applies a local-in-time decomposition on each interval $t_{n},t_{n+1}]$, combined with an EWI for discretizing the resulting subproblems. For a fixed $n\geq 0$, the data at $t=t_{n}$ are assumed to satisfy
(30) $u(\mathbf{x},t_{n})=\phi_{1}^{n}(\mathbf{x})=O(1),\qquad\partial_{t}u(\mathbf{x},t_{n})=\varepsilon^{-2}\phi_{2}^{n}(\mathbf{x})=O(\varepsilon^{-2}),\qquad\mathbf{x}\in\mathbb{R}^{d},$
To derive a decomposition valid for the full regime $\varepsilon\in(0,1]$, we introduce a WKB expansion with a remainder by writing $u(\mathbf{x},t)=u(\mathbf{x},t_{n}+s)$ in the form
(31) $u(\mathbf{x},t_{n}+s)=e^{is/\varepsilon^{2}}z^{n}(\mathbf{x},s)+e^{-is/\varepsilon^{2}}\,\overline{z^{n}}(\mathbf{x},s)+r^{n}(\mathbf{x},s),\quad\mathbf{x}\in\mathbb{R}^{d},\quad 0\leq s\leq\tau.$
This yields a multiscale decomposition with the $\varepsilon$-frequency for the NKGE [bao2014mti, bao2013ewi].
(32a) $2i\partial_{s}z^{n}(\mathbf{x},s)+\varepsilon^{2}\partial_{ss}z^{n}(\mathbf{x},s)-\Delta z^{n}(\mathbf{x},s)+3\lambda|z^{n}(\mathbf{x},s)|^{2}z^{n}(\mathbf{x},s)=0,$
(32b) $\varepsilon^{2}\partial_{ss}r^{n}(\mathbf{x},s)-\Delta r^{n}(\mathbf{x},s)+\frac{1}{\varepsilon^{2}}r^{n}(\mathbf{x},s)+f_{r}(z^{n}(\mathbf{x},s),r^{n}(\mathbf{x},s);s)=0,$
with the well-prepared initial data for $z^{n}$ and small initial data for $r^{n}$ as [bao2014mti, bao2019comp]
(33) $\left\{\begin{split}&z^{n}(\mathbf{x},0)=\frac{1}{2}\left[\phi_{1}^{n}(\mathbf{x})-i\phi_{2}^{n}(\mathbf{x})\right],\\ &\partial_{s}z^{n}(x,0)=\frac{i}{2}\left[-\Delta z^{n}(\mathbf{x},0)+3\lambda|z^{n}(\mathbf{x},0)|^{2}z^{n}(\mathbf{x},0)\right],\\ &r^{n}(\mathbf{x},0)=0,\qquad\partial_{s}r^{n}(\mathbf{x},0)=-\partial_{s}z^{n}(\mathbf{x},0)-\partial_{s}\overline{z^{n}}(\mathbf{x},0),\end{split}\right$.
where
$\begin{split}f_{r}\left(z,r;s\right)=&\lambda e^{3is/\varepsilon^{2}}z^{3}+\lambda e^{-3is/\varepsilon^{2}}\overline{z}^{3}+3\lambda\left(e^{2is/\varepsilon^{2}}z^{2}+e^{-2is/\varepsilon^{2}}\overline{z}^{2}\right)r\\ &+3\lambda\left(e^{is/\varepsilon^{2}}z+e^{-is/\varepsilon^{2}}\overline{z}\right)r^{2}+6\lambda|z|^{2}r+\lambda r^{3}.\end{split}$
3.2 Neural multiscale decomposition (NeuralMD)
However, the MTI method is limited by its reliance on local-in-time solution behavior. In the nonrelativistic regime, it captures only the limiting dynamics on local time scales, which prevents it from effectively resolving long-time multiscale frequency interactions and achieving global-in-time convergence. When long-time behavior is inferred through the accumulation of local solutions, errors may propagate and amplify. In contrast, PINNs employ randomized collocation over the full time domain, enabling the optimization of the solution space globally in time. By enforcing the initial condition, boundary conditions, and governing equations across the entire temporal interval, PINNs provide a more accurate description of the global solution dynamics.
Based on this, we combine the multiscale decomposition of the MTI method with the global-time optimization framework of PINNs to develop an adaptive dropping time oscillation strategy, referred to as NeuralMD. The NeuralMD method follows a two-stage pretraining framework (see Fig. 7). The first stage solves the NLSW (32a) with well-prepared initial data, enabling efficient removal of high-frequency time oscillation, particularly in the nonrelativistic regime. The second stage focuses on the remainder equation (32b), whose initial data are small but exhibit high-frequency behavior, and provides an effective compensation of the remainder amplitude in regions with slight time oscillation.
Figure 7: Overall architecture of NeuralMD.The single input point is perturbed in time to form point sets across multiscale regions. The model gradients are propagated backward across multiple regions.
In this work, we employ a global-in-time multiscale decomposition of the NKGE solution of the form
(34) $u(\mathbf{x},t)=e^{it/\varepsilon^{2}}z(\mathbf{x},t)+e^{-it/\varepsilon^{2}}\,\overline{z}(\mathbf{x},t)+r(\mathbf{x},t),\quad x\in\mathbb{R}^{d},\quad 0\leq t\leq T.$
Pre-training stage I
The first stage focuses on the modulated amplitude equation (32a) arising from the multiscale frequency decomposition,
(35) $2i\,\partial_{t}z(x,t)+\varepsilon^{2}\partial_{tt}z(x,t)-\Delta z(x,t)+3\lambda|z(x,t)|^{2}z(x,t)=0,$
with initial conditions
$z(x,0)=\frac{1}{2}\bigl[\phi_{1}(x)-i\phi_{2}(x)\bigr],\quad\partial_{t}z(x,0)=\frac{i}{2}\Bigl[-\Delta z(x,0)+3\lambda|z(x,0)|^{2}z(x,0)\Bigr].$
Let $z_{\theta}=z_{\theta}^{\mathrm{re}}+i\,z_{\theta}^{\mathrm{im}}$ be the neural network approximation. The residual loss associated with (35) is given by
$\mathcal{L}_{\mathrm{Res}}^{z}(\theta_{1})=$ $\mathbb{E}_{(x,t)\in P_{f}}\bigl\|2\,\partial_{t}z_{\theta}^{\mathrm{re}}+\varepsilon^{2}\partial_{tt}z_{\theta}^{\mathrm{im}}-\Delta z_{\theta}^{\mathrm{im}}+3\lambda|z_{\theta}|^{2}z_{\theta}^{\mathrm{im}}\bigr\|_{L^{2}}^{2}$
(36) $+\mathbb{E}_{(x,t)\in P_{f}}\bigl\|-2\,\partial_{t}z_{\theta}^{\mathrm{im}}+\varepsilon^{2}\partial_{tt}z_{\theta}^{\mathrm{re}}-\Delta z_{\theta}^{\mathrm{re}}+3\lambda|z_{\theta}|^{2}z_{\theta}^{\mathrm{re}}\bigr\|_{L^{2}}^{2}.$
The initial-condition loss is given by
$\mathcal{L}_{\mathrm{Ic}}^{z}(\theta_{1})=$ $\mathbb{E}_{x\in P_{0}}\left(\frac{1}{2}\|z_{\theta}^{\mathrm{re}}(x,0)-\phi_{1}(x)\|_{L^{2}}^{2}+\frac{1}{2}\|\Delta z_{\theta}^{\mathrm{im}}(x,0)-3\lambda|z_{\theta}|^{2}z_{\theta}^{\mathrm{im}}(x,0)\|_{L^{2}}^{2}\right)$
(37) $+\mathbb{E}_{x\in P_{0}}\left(\frac{1}{2}\|z_{\theta}^{\mathrm{im}}(x,0)-\phi_{2}(x)\|_{L^{2}}^{2}+\frac{1}{2}\|-\Delta z_{\theta}^{\mathrm{re}}(x,0)+3\lambda|z_{\theta}|^{2}z_{\theta}^{\mathrm{re}}(x,0)\|_{L^{2}}^{2}\right).$
And the period boundary loss is given by
$\mathcal{L}^{z}_{\mathrm{Bd}}(\theta_{1})$ $=\mathbb{E}_{(\mathbf{x},t)\in P_{b}}\left(\left\|z^{\mathrm{re}}_{\theta_{1}}(a,t_{a})-z^{\mathrm{re}}_{\theta_{1}}(b,t_{b})\right\|_{L^{2}}^{2}+\left\|\partial_{x}z^{\mathrm{re}}_{\theta_{1}}(a,t_{a})-\partial_{x}z^{\mathrm{re}}_{\theta_{1}}(b,t_{b})\right\|_{L^{2}}^{2}\right)$
(38) $+\mathbb{E}_{(\mathbf{x},t)\in P_{b}}\left(\left\|z^{\mathrm{im}}_{\theta_{1}}(a,t_{a})-z^{\mathrm{im}}_{\theta_{1}}(b,t_{b})\right\|_{L^{2}}^{2}+\left\|\partial_{x}z^{\mathrm{im}}_{\theta_{1}}(a,t_{a})-\partial_{x}z^{\mathrm{im}}_{\theta_{1}}(b,t_{b})\right\|_{L^{2}}^{2}\right),$
The total loss for the first stage is then
(39) $\mathcal{L}^{z}(\theta_{1})=\lambda_{\mathrm{Res}}\mathcal{L}_{\mathrm{Res}}^{z}(\theta_{1})+\lambda_{\mathrm{Ic}}\mathcal{L}_{\mathrm{Ic}}^{z}(\theta_{1})+\lambda_{\mathrm{Bd}}\mathcal{L}_{\mathrm{Bd}}^{z}(\theta_{1}).$
In the multiscale decomposition (31), the rapidly oscillatory factors $e^{\pm it/\varepsilon^{2}}$ are extracted explicitly from the solution $u$, and the slowly varying envelope $z(\mathbf{x},t)$ evolves on the slow time scale $O(1)$ according to (32a). Consequently, the temporal wavelength of $z$ is $O(1)$ and its temporal spectrum is localized in a band $|k|=O(1)$ that is independent of $\varepsilon$. Taking the temporal Fourier transform of (32a) with respect to $s$ (and denoting by $\xi$ the spatial frequency), the linear part of the residual for $z^{n}$ in the frequency domain carries the symbol
(40) $\sigma_{z}(k,\xi)=-2k-\varepsilon^{2}k^{2}+|\xi|^{2},$
so that the corresponding coefficient satisfies
(41) $\bigl|\sigma_{z}(k,\xi)\bigr|=\bigl|-2k-\varepsilon^{2}k^{2}+|\xi|^{2}\bigr|=O(1),\qquad|k|=O(1),$
uniformly in $\varepsilon\in(0,1]$ for the physically relevant band of temporal frequencies.
When training a PINN on $z$ with a residual-based physics loss associated with (32a), the dominant contribution of temporal frequency $k$ to the gradient with respect to a parameter $\theta_{j}$ can be written in analogy with the NKGE case as
(42) $\frac{\partial\mathcal{L}^{z}(\theta_{1},k)}{\partial\theta_{1j}}\approx\bigl|\sigma_{z}(k,\xi)\bigr|\,\bigl\lVert z_{\theta_{1}}(k,\xi)-z(k,\xi)\bigr\rVert\exp\!\left(-\frac{\pi|k|}{2|w_{j}|}\right)\,F_{j}.$
where $F_{j}$ is an $O(1)$ factor incorporating the contribution of the nonlinear term and spatial derivatives on this frequency band. Since the temporal spectrum of $z$ is concentrated at $|k|=O(1)$, the coefficient $\bigl|\sigma_{z}(k,\xi)\bigr|$ remains bounded and does not grow like $O(\varepsilon^{-2})$ as $\varepsilon\to 0$.
At the same time, the network sensitivity factor $\exp\!\left(-\pi|k|/(2|w_{j}|)\right)$ only attenuates genuinely high temporal frequencies, whereas the relevant frequencies for $z$ stay in the low-frequency regime where this factor is $O(1)$. Therefore, in the multiscale decomposition, the PINN is trained on a slowly varying envelope $z$ whose temporal frequencies are compatible with the intrinsic spectral bias of the network, and the mechanism leading to the breakdown of spectral bias for the original NKGE in the nonrelativistic limit no longer occurs.
Pre-training stage II
Then, in the second stage, the network prediction $z_{\theta}$ from the first stage is used to decouple the remainder equation. The remainder $r(x,t)$ satisfies, for $t>0$,
(43) $\varepsilon^{2}\partial_{tt}r(\mathbf{x},t)-\Delta r(\mathbf{x},t)+\varepsilon^{-2}r(\mathbf{x},t)+f_{r}\bigl(z(\mathbf{x},t),r(\mathbf{x},t);t\bigr)=0,$
with initial conditions
(44) $r(\mathbf{x},0)=0,\qquad\partial_{t}r(\mathbf{x},0)=-\partial_{t}z(\mathbf{x},0)-\partial_{t}\overline{z}(\mathbf{x},0).$
Let $r_{\theta_{2}}$ be the neural network approximation to the remainder. The three loss components are given by
(45) $\left\{\begin{aligned} \mathcal{L}_{\mathrm{Res}}^{r}(\theta_{2})&=\mathbb{E}_{(\mathbf{x},t)\in P_{f}}\Bigl\|\varepsilon^{2}\partial_{tt}r_{\theta_{2}}-\Delta r_{\theta_{2}}+\varepsilon^{-2}r_{\theta_{2}}+f\bigl(z_{\theta_{1}},r_{\theta_{2}}\bigr)\Bigr\|_{L^{2}}^{2},\\[4.0pt] \mathcal{L}_{\mathrm{Ic}}^{r}(\theta_{2})&=\mathbb{E}_{\mathbf{x}\in P_{0}}\left(\|r_{\theta_{2}}(\mathbf{x},0)\|_{L^{2}}^{2}+\|\partial_{t}r_{\theta_{2}}(\mathbf{x},0)+(\partial_{t}z_{\theta_{1}}(\mathbf{x},0)+\mathrm{c.c.})\|_{L^{2}}^{2}\right),\\[4.0pt] \mathcal{L}^{r}_{\mathrm{Bd}}(\theta_{2})&=\mathbb{E}_{(\mathbf{x},t)\in P_{b}}\left(\left\|r_{\theta_{2}}(a,t_{a})-r_{\theta_{2}}(b,t_{b})\right\|_{L^{2}}^{2}+\left\|\partial_{x}r_{\theta_{2}}(a,t_{a})-\partial_{x}r_{\theta_{2}}(b,t_{b})\right\|_{L^{2}}^{2}\right).\end{aligned}\right$.
The total loss for the second stage is
(46) $\mathcal{L}^{r}(\theta_{2})=\lambda_{\mathrm{Res}}\mathcal{L}_{\mathrm{Res}}^{r}(\theta_{2})+\lambda_{\mathrm{Ic}}\mathcal{L}_{\mathrm{Ic}}^{r}(\theta_{2})+\lambda_{\mathrm{Bd}}\mathcal{L}_{\mathrm{Bd}}^{r}(\theta_{2}),$
To enforce the initial conditions automatically, we parameterize the outputs as
$r_{\theta_{2}}(\mathbf{x},t)=(1+t)e^{-t}\,r_{0}(\mathbf{x})+te^{-t}\,\partial_{t}r_{0}(\mathbf{x})+\bigl(1-e^{-t}-te^{-t}\bigr)\,r_{\theta_{2}}(\mathbf{x},t).$
where $r_{0}$ and $\partial_{t}r_{0}$ denote the prescribed initial data.
In contrast to the slowly varying envelope $z_{\theta_{1}}$, the remainder $r_{\theta_{2}}$ in the multiscale decomposition still contains oscillatory components at the carrier frequencies $\pm\varepsilon^{-2}$ through the nonlinear interaction terms appearing in $f_{r_{\theta_{2}}}$. Although the initial data for $r_{\theta_{2}}$ in (33) are well prepared and satisfy (44). so that $r_{\theta_{2}}$ remains small in appropriate norms uniformly for $\varepsilon\in(0,1]$, the temporal frequencies present in $r_{\theta_{2}}$ are still of order $k\sim\varepsilon^{-2}$. This follows from the oscillatory factors $e^{\pm it/\varepsilon^{2}},e^{\pm 2it/\varepsilon^{2}},e^{\pm 3it/\varepsilon^{2}}$ contained in $f_{r_{\theta_{2}}}$, which excite the high-frequency temporal modes inherited from the original NKGE.
Taking the temporal Fourier transform of the remainder equation (32b) yields the linear frequency-domain coefficient
(47) $\sigma_{r}(k,\xi)=-\varepsilon^{2}k^{2}+|\xi|^{2}+\varepsilon^{-2},$
so that in the oscillatory band $k\sim\varepsilon^{-2}$ one has
(48) $\bigl|\sigma_{r}(k,\xi)\bigr|=\bigl|-\varepsilon^{2}k^{2}+|\xi|^{2}+\varepsilon^{-2}\bigr|=O(\varepsilon^{-2}),\qquad k\sim\varepsilon^{-2},$
Hence, the residual associated with $r^{n}$ assigns large nominal weights to these high temporal frequencies, in the same manner as in the original NKGE.
If a PINN is used to approximate $r^{n}$, the dominant contribution of frequency $k$ to the gradient with respect to a trainable parameter $\theta_{j}$ satisfies
(49) $\frac{\partial\mathcal{L}^{r}(\theta_{2},k)}{\partial\theta_{j}}\approx\bigl|\sigma_{r}(k,\xi)\bigr|\,\bigl\lVert r_{\theta_{2}}(k,\xi)-r(k,\xi)\bigr\rVert\exp\!\left(-\frac{\pi|k|}{2|w_{j}|}\right)\,G_{j}.$
where $G_{j}=O(1)$ depends on $\theta_{j}$ and the spatial derivatives.
Since $|k|\sim\varepsilon^{-2}$ in the active oscillatory band, the sensitivity factor $\exp(\pi|k|/(2|w_{j}|))$ is exponentially small during the early stages of training when $|w_{j}|\ll 1$. Therefore, even though $r^{n}$ itself has small amplitude, its high-frequency temporal content induces a secondary form of spectral bias: the residual loss allocates large gradients to high frequencies, but the network response at these frequencies is exponentially suppressed. This mismatch leads to slow convergence and high-frequency drift in the training dynamics of the $r^{n}$-component, mirroring (at a reduced scale) the breakdown of spectral bias observed in the original NKGE.
Error criterion
After the two-stage pretraining, using the pretrained weights from the amplitude and remainder networks, we obtain the predictions $z_{\theta_{1}^{\ast}}$ and $r_{\theta_{2}^{\ast}}$ and reconstruct the oscillation solution by
(50) $\left\{\begin{aligned} u_{\theta_{1}^{\ast}}(\mathbf{x},t)&=e^{it/\varepsilon^{2}}z_{\theta_{1}^{\ast}}(\mathbf{x},t)+\mathrm{c.c.},\\ u_{(\theta_{1}^{\ast},\theta_{2}^{\ast})}(\mathbf{x},t)&=e^{it/\varepsilon^{2}}z_{\theta_{1}^{\ast}}(\mathbf{x},t)+\mathrm{c.c.}+r_{\theta_{2}^{\ast}}(\mathbf{x},t),\end{aligned}\right$.
Further, we introduce an error-based criterion to determine whether the remainder should be incorporated for amplitude compensation, as
(51) $u_{\theta^{\ast}}\leftarrow\mathrm{min}\{\|u_{\theta_{1}^{\ast}}-\hat{u}\|,\|u_{(\theta_{1}^{\ast},\theta_{2}^{\ast})}-\hat{u}\|\}$.
where, $\hat{u}$ is the ground truth solution of NKGE.
3.3 Gated gradient correlation correction
In this section, we investigate the impact of multiscale frequency decomposition on the propagation failure mode induced by PINNs. For the amplitude equation with mild-frequency time oscillation, the time angular frequency spectrum density is given by
(52) $S_{x}(k,\varepsilon)\sim\varepsilon f(k\varepsilon).$
The gradient decorrelation is expressed as
(53) $\langle k^{2}\rangle_{S,x}=\frac{\int_{0}^{\infty}k^{2}S_{x}(k,\varepsilon)\,dk}{\int_{0}^{\infty}S_{x}(k,\varepsilon)\,dk}\sim\varepsilon^{2}\int_{0}^{\infty}k^{2}f(k\varepsilon)\,dk.$
The gradient correlation decay satisfies
(54) $G_{z_{\theta_{1}}}(x,\delta t)\approx A(x)-\frac{1}{2}A(x)\left(\varepsilon^{2}\int_{0}^{\infty}k^{2}f(k\varepsilon)\,dk\right)\delta t^{2}+O(\delta t^{4}).$
As $\varepsilon\to 0$, $\langle k^{2}\rangle_{S,x}$ decreases gradually, and when $\varepsilon\to 0$, $\langle k^{2}\rangle_{S,x}\to 0$, indicating that the gradient correlation between adjacent time steps remains constant. This is consistent with the convergence of the semi-classical limit equation to the limiting equation as $\varepsilon\to 0$.
The semi-classical limit equation
(55) $2i\partial_{t}z(x,t)+\varepsilon^{2}\partial_{tt}z(x,t)-\Delta z(x,t)+3\lambda|z(x,t)|^{2}z(x,t)=0$
has a negligible effect on the gradient correlation. The limiting equation
(56) $2i\partial_{t}z(x,t)-\Delta z(x,t)+3\lambda|z(x,t)|^{2}z(x,t)=0$
is unaffected by gradient correlation. Therefore, solving the amplitude equation with PINNs does not induce propagation failure.
For the remainder equation, although the initial data are small, its temporal oscillatory characteristics remain consistent with the original equation. The time angular frequency spectrum density is given by
(57) $S_{x}(k,\varepsilon)\sim\frac{1}{\varepsilon}f\left(\frac{k}{\varepsilon}\right).$
The gradient decorrelation is then
(58) $\langle k^{2}\rangle_{S,x}\sim\frac{1}{\varepsilon^{2}}\int_{0}^{\infty}k^{2}f\left(\frac{k}{\varepsilon}\right)\,dk.$
The gradient correlation is expressed as
(59) $G_{r_{\theta_{2}}}(x,\delta t)\approx A(x)-\frac{1}{2}A(x)\left(\frac{1}{\varepsilon^{2}}\int_{0}^{\infty}k^{2}f\left(\frac{k}{\varepsilon}\right)\,dk\right)\delta t^{2}+O(\delta t^{4}).$
As $\varepsilon$ decreases, $\langle k^{2}\rangle_{S,x}$ increases, and the gradient correlation between adjacent time steps decays more rapidly. When $\varepsilon\to 0$, the gradient correlation decays sharply, and PINNs solving the remainder equation induce propagation failure. This results in the failure of the second stage pretraining of NeuralMD, where the remainder is discarded based on the error criterion. However, when $\varepsilon$ is slightly reduced from 1, PINNs’ propagation failure is induced during long-time predictions, and NeuralMD prematurely discards the remainder. To address this, we aim to enhance PINNs’ ability to mitigate propagation failure and shift the failure boundary towards the transition region.
To mitigate the propagation failure mode, we incorporate temporal causality into NeuralMD to prevent premature convergence to a trivial solution (low residual, high error), which halts correct solution propagation. We address this from two perspectives: from the model perspective, we apply gradient correlation correction by enhancing the inner product of parameter gradients at adjacent spatiotemporal points, increasing the impact of updates on neighboring points, and ensuring continuous oscillatory evolution. From the sampling perspective, we introduce time causality gating to prioritize high-residual fronts within the current time window. Once the error in this window converges, we progressively advance to future time steps, improving the performance of gradient correlation correction.
3.3.1 Gated gradient flow
The gradient correlation in (15) ensures uniform propagation of gradients over the entire time span, but it does not establish causal relationships between gradients across time. As a result, precise gradient information from earlier times is difficult to propagate to the future, and gradients at later times are often amplified incorrectly. To address this, we propose a gate gradient flow strategy that introduces dynamic characteristics to gradient correlation.
The core idea is to use a time-dependent gate function $h(t)$ that explicitly enforces causality by revealing only a portion of the time domain to NeuralMD during training. Specifically, we define the continuous gate function as
(60) $h(t)=\frac{1-\tanh(\alpha(\tilde{t}-\gamma))}{2},$
where $\gamma$ is the shift parameter controlling the fraction of time revealed to the model, $\alpha=5$ controls the steepness of the gate, and $\tilde{t}=t/T$ is the normalized time.
Taking the amplitude equation (25) as an example, we use $h(t)$ to modify (15) and obtain a gated gradient correlation metric, as
[…]
Compared to (15), (61) evolves from a static similarity measure to a dynamic object that changes throughout the training process. In the early training, the model relies more on the strong correlation of gradients from earlier times, while as the causal gate opens, later time gradients gradually contribute to the optimization. The spectral structure of gradient correlation changes with the gate, altering the network’s sensitivity to different time regions.
The shift parameter $\gamma$ is updated at each iteration according to the following rule
(62) $\gamma^{i+1}=\gamma^{i}+\eta e^{-\epsilon G_{z_{\theta_{1}}}},$
where $\eta$ is the learning rate, $\epsilon$ is a tolerance parameter that controls when the gate shifts, and $i$ is the iteration number. With $\eta=1\text{e-}3$, if the expected value of $e^{-\epsilon G_{z_{\theta_{1}}}}$ over 1000 iterations is 0.1, $\gamma$ will change by 0.1 after 1000 iterations.
For the tanh gate function, $\gamma$ typically ranges from 0.5-0.5 to $1.5$. When the gradient correlation $G_{z_{\theta_{1}}}$ approaches zero, the update magnitude $e^{-\epsilon G_{z_{\theta_{1}}}}$ approaches 1, which can cause abrupt changes in the gate. To mitigate this, we apply a magnitude clipping scheme, similar to gradient clipping, as follows
(63) $\gamma^{i+1}=\gamma^{i}+\eta\min(e^{-\epsilon G_{z_{\theta_{1}}}},\Delta_{max}).$
where $\Delta_{max}$ is the maximum allowed update magnitude, typically set to $\Delta_{max}=0.1$. The choice of $\Delta_{max}$ depends on $\eta$.
The gate function $h$ is not restricted to the tanh function. Any continuous, monotonically decreasing function $h$ dependent on a shift parameter $\gamma$, where increasing $\gamma$ raises the gate value, can serve as a valid gate. An alternative gate funcion is a composition of ReLU and $\tanh$, given by $h=\text{ReLU}(-\tanh(\alpha(t-\gamma)))$, as shown in Fig. 8. This gate function enforces a stricter threshold, pushing the gate value to 0 after a certain time cutoff.
Figure 8: Gate function by Relu and tanh.
3.3.2 Random time perturbation
Given a single collocation point $(\mathbf{x},t)\in\Omega\subseteq\mathbb{R}^{d+1}$ with temporal coordinate $t$ and spatial coordinate $\mathbf{x}\in\mathbb{R}^{d}$, we augment it by perturbing only its time coordinate within multiscale temporal regions. In contrast to PINNsFormer [zhao2023former], which detaches gradients during augmentation, and ProPINNs [wu2025propinn], where perturbing high-dimensional physical regions is unsuitable for handling time oscillation, NeuralMD treats time perturbation as a differentiable layer.
Importantly, the temporal perturbation is designed to explicitly align with the gated gradient flow behavior ((61)) of the amplitude and remainder equations under the influence of the gate function $h(t)$. For the amplitude equation, whose spectrum
(64) $S_{x}(k,\varepsilon)\sim\varepsilon f(k\varepsilon),$
implies a slowly decaying gated gradient correlation
(65) $G_{z_{\theta_{1}}}(t,t+\delta t)\approx A(t)-\tfrac{1}{2}A(t)\langle k^{2}\rangle_{S,t}\,\delta t^{2}+O(\delta t^{4})\quad\text{with}\quad\langle k^{2}\rangle_{S,t}\sim\varepsilon^{2}\to 0,$
the perturbation preserves stable gradient behavior while respecting causality enforced by $h(t)$. For the remainder equation, whose oscillatory spectrum
(66) $S_{x}(k,\varepsilon)\sim\tfrac{1}{\varepsilon}f(k/\varepsilon),$
induces rapid correlation decay
(67) $G_{r_{\theta_{2}}}(t,t+\delta t)\approx A(t)-\tfrac{1}{2}A(t)\langle k^{2}\rangle_{S,t}\,\delta t^{2}+O(\delta t^{4}).\qquad\langle k^{2}\rangle_{S,t}\sim\tfrac{1}{\varepsilon^{2}}\to\infty.$
the perturbation provides a mechanism to correct gradient decorrelation by aggregating gradients across controlled temporal neighborhoods, modulated by the dynamic gate $h(t)$.
We now introduce the multiscale time perturbation formally, as
(68) $\left\{\begin{aligned} &\operatorname{Diff\text{-}Aug}(\mathbf{x},t)=\bigg\{(\mathbf{x},t),\Big\{\{(\mathbf{x},t+\delta t_{l}^{\,i})\}_{i=1}^{k_{l}}\Big\}_{l=1}^{\text{\#scale}}\bigg\},\\[5.69054pt] &\mathbf{X}_{\text{point}}=\mathcal{P}(\mathbf{x},t),\qquad\left\{\mathbf{X}_{\text{region}}^{i,l}\right\}_{i=1}^{k_{l}}=\left\{\mathcal{P}(\mathbf{x},t+\delta t_{l}^{\,i})\right\}_{i=1}^{k_{l}}.\end{aligned}\right$.
Here $\{\delta t_{l}^{\,i}\}_{i=1}^{k_{l}}$ are random temporal perturbations belonging to the $r$-th time region of size $R_{l},R_{l}]$, and $k_{l}$ denotes the number of perturbations per scale. The number of temporal scales is denoted as #scale. All representations $\mathbf{X}_{\text{point}},\mathbf{X}_{\text{region}}^{i,l}\in\mathbb{R}^{d_{\text{model}}}$ are obtained via a lightweight coordinate encoder
(69) $\mathcal{P}:\mathbb{R}^{d+1}\to\mathbb{R}^{d_{\text{model}}}.$
The use of multiscale temporal regions simultaneously (i) captures the intrinsic multiscale dynamics of the NKGE system, and (ii) resembles adaptive temporal discretizations where each time point selectively interacts with multiple temporal neighborhoods, gated by $h(t)$ to enforce causality.
From the forward-model view, multiscale perturbation enlarges the temporal receptive field. From the backward view, it aggregates gradients from neighboring perturbed points under the influence of the gate function $h(t)$,
(70) $\left\{\begin{aligned} \nabla_{\theta^{\mathcal{P}}_{1}}\mathcal{L}^{z}(\theta_{1})\;\leftarrow\;\sum_{l=1}^{\#\text{scale}}\sum_{i=1}^{k_{l}}\sqrt{h(t)}\sqrt{h(t+\delta t_{l}^{\,i})}\left\langle\left.\frac{\partial z_{\theta_{1}}}{\partial\theta^{\mathcal{P}}_{1}}\right|_{t},\;\left.\frac{\partial z_{\theta_{1}}}{\partial\theta^{\mathcal{P}}_{1}}\right|_{t+\delta t_{l}^{\,i}}\right\rangle,\\ \nabla_{\theta^{\mathcal{P}}_{2}}\mathcal{L}^{r}(\theta_{2})\;\leftarrow\;\sum_{l=1}^{\#\text{scale}}\sum_{i=1}^{k_{l}}\sqrt{h(t)}\sqrt{h(t+\delta t_{l}^{\,i})}\left\langle\left.\frac{\partial r_{\theta_{2}}}{\partial\theta^{\mathcal{P}}_{2}}\right|_{t},\;\left.\frac{\partial r_{\theta_{2}}}{\partial\theta^{\mathcal{P}}_{2}}\right|_{t+\delta t_{l}^{\,i}}\right\rangle,\end{aligned}\right$.
which directly enhances the gated gradient correlation, by
(71) $\left\{\begin{aligned} G_{z_{\theta_{1}}}(t,\delta t_{l}^{\,i})=\sqrt{h(t)}\sqrt{h(t+\delta t_{l}^{\,i})}\left\langle\left.\frac{\partial z_{\theta_{1}}}{\partial\theta^{\mathcal{P}}_{1}}\right|_{t},\;\left.\frac{\partial z_{\theta_{1}}}{\partial\theta^{\mathcal{P}}_{1}}\right|_{t+\delta t_{l}^{\,i}}\right\rangle,\\ G_{r_{\theta_{2}}}(t,\delta t_{l}^{\,i})=\sqrt{h(t)}\sqrt{h(t+\delta t_{l}^{\,i})}\left\langle\left.\frac{\partial r_{\theta_{2}}}{\partial\theta^{\mathcal{P}}_{2}}\right|_{t},\;\left.\frac{\partial r_{\theta_{2}}}{\partial\theta^{\mathcal{P}}_{2}}\right|_{t+\delta t_{l}^{\,i}}\right\rangle.\end{aligned}\right$.
where, the shift parameter $\gamma$ in $h(t)$ is updated adaptively based on the average $G_{z_{\theta_{1}}}$ over iterations, ensuring perturbations only activate in revealed time windows.
For the amplitude equation, where $G_{z_{\theta_{1}}}$ and $G_{r_{\theta_{2}}}$ naturally decay slowly, the gated perturbation preserves stable propagation. For the remainder equation, where $G_{r_{\theta_{2}}}$ suffers rapid correlation decay (propagation failure), the gated perturbation provides a controlled multiscale correction, effectively slowing the decay and stabilizing long-time training dynamics.
3.3.3 Multiscale time mixing
After the shared projection layer, each collocation point generates $(1+\sum_{l=1}^{\text{\#scale}}k_{l})$ representations corresponding to its multiscale temporal perturbations. Unlike existing methods to model complex spatiotemporal dependencies, our setting focuses exclusively on multiscale time mixing. Attention requires pairwise inner products between representations, leading to prohibitively large forward and backward computational costs, especially since PINNs must compute high-order gradients.
Instead of modeling all mutual dependencies, we introduce an efficient multiscale time mixing mechanism that aggregates representations within each time region and mixes them across different temporal scales. This is consistent with the multiscale decomposition of the NKGE system:
  • The amplitude equation has a low-frequency spectrum $S_{x}(k,\varepsilon)\sim\varepsilon f(k\varepsilon)$, producing slowly varying, highly correlated temporal features.
  • The remainder equation has a high-frequency spectrum $S_{x}(k,\varepsilon)\sim\varepsilon^{-1}f(k/\varepsilon)$, producing rapidly decorrelating temporal features.
Pooling over temporal regions thus produces a hierarchy of feature components aligned with the multiscale temporal structures imposed by the PDE: from low-oscillatory (amplitude-like) to high-oscillatory (remainder-like) regimes. Because all perturbed points follow the same operator, but correspond to different effective temporal resolutions, the resulting representations form a family of PDEs with varying coefficients [graham2007]. The relations among these coefficient-modified PDEs are much more stable than the relations among raw collocation points, enabling a simple linear mixing layer without the need for attention.
Formally, for each temporal scale $l$, we first pool the representations from perturbations $\{t+\delta t_{l}^{\,i}\}_{i=1}^{k_{l}}$ with gate function $h(t)$, as
(72) $\mathbf{X}_{\text{region}}^{l}$ $=\operatorname{Pooling}\Big(\{\mathbf{X}_{\text{region}}^{i,l}\}_{i=1}^{k_{l}}\cdot h(t+\delta t_{l}^{\,i})\Big),\qquad l=1,\ldots,\text{\#scale},$
$\mathbf{X}$ $=\mathcal{M}\!\left(\mathbf{X}_{\text{point}},\mathbf{X}_{\text{region}}^{1},\ldots,\mathbf{X}_{\text{region}}^{\text{\#scale}}\right),$
where $\mathcal{M}:\mathbb{R}^{(1+\text{\#scale})\times d_{\text{model}}}\rightarrow\mathbb{R}^{d_{\text{model}}}$ is a lightweight MLP that mixes multiscale temporal features. The mixed representation $\mathbf{X}$ is then passed through another decoder MLP $\mathcal{H}:\mathbb{R}^{d_{\text{model}}}\rightarrow\mathbb{R}^{m}$, which produces the predicted solution
(73) $z_{\theta_{1}}(\mathbf{x},t)=\mathcal{H}(\mathbf{X})\in\mathbb{R}^{m},\quad r_{\theta_{2}}(\mathbf{x},t)=\mathcal{H}(\mathbf{X})\in\mathbb{R}^{m}.$
This multiscale time mixing mechanism effectively combines Low-frequency, highly correlated components (reflecting the amplitude equation) with high-frequency, rapidly decorrelating components (reflecting the remainder equation). Thus, it enhances NeuralMD’s ability to maintain time gradient correlation and mitigate propagation failure during long-time predictions.
3.3.4 Justification of gradient correlation
{assumption}
[Gradient correlation across time regions] For the NeuralMD method with $z_{\theta_{1}}$ and $r_{\theta_{2}}$, we assume the existence of a time region size $R>0$ such that for all $0\leq t\leq T$ and $0\leq t+\delta t\leq T$ with $\|\delta t\|\leq R$, the following conditions hold
$\left\langle\left.\frac{\partial z_{\theta_{1}}}{\partial\theta}\right|_{t},\left.\frac{\partial z_{\theta_{1}}}{\partial\theta}\right|_{t+\delta t}\right\rangle\geq 0,\quad\left\langle\left.\frac{\partial r_{\theta_{2}}}{\partial\theta_{2}}\right|_{t},\left.\frac{\partial r_{\theta_{2}}}{\partial\theta_{2}}\right|_{t+\delta t}\right\rangle\geq 0.$
Proof.
We first note that, similarly to Assumption 3.3.4, if we relax the constraint on the time region size $R$ from “positive” to “non-negative”, Assumption 3.3.4 is trivially satisfied. Indeed, taking $R=0$ yields
(74) $\left\langle\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t},\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t}\right\rangle=\bigg\|\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t}\bigg\|^{2}\geq 0,$
and
(75) $\left\langle\left.\frac{\partial r_{\theta_{2}}}{\partial\theta_{2}}\right|_{t},\left.\frac{\partial r_{\theta_{2}}}{\partial\theta_{2}}\right|_{t}\right\rangle=\bigg\|\left.\frac{\partial r_{\theta_{2}}}{\partial\theta_{2}}\right|_{t}\bigg\|^{2}\geq 0,$
for all $t\in[0,T]$.
In the following we keep the requirement that $R>0$. For brevity, define
(76) $g_{1}(t):=\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t},\qquad g_{2}(t):=\left.\frac{\partial r_{\theta_{2}}}{\partial\theta_{2}}\right|_{t}.$
Assume that the mixed derivative $\partial^{2}z_{\theta_{1}}/(\partial\theta_{1}\,\partial t)$ is bounded on $[0,T]$, i.e., there exists $L_{1}>0$ such that
(77) $\bigg\|\frac{dg_{1}}{dt}(s)\bigg\|\leq L_{1},\quad\forall s\in[0,T].$
Then $g_{1}$ is $L_{1}$-Lipschitz in $t$, and for any $t,t+\delta t\in[0,T]$ we have
(78) $\|g_{1}(t+\delta t)-g_{1}(t)\|\leq L_{1}|\delta t|.$
Fix any $t\in[0,T]$.
  • If $\|g_{1}(t)\|\neq 0$, then
    (79) $\langle g_{1}(t),g_{1}(t)\rangle=\|g_{1}(t)\|^{2}>0,$
    For any $\delta t$ with $t,t+\delta t\in[0,T]$, we estimate
    (80) $\langle g_{1}(t),g_{1}(t+\delta t)\rangle$ $=\langle g_{1}(t),g_{1}(t)\rangle+\langle g_{1}(t),g_{1}(t+\delta t)-g_{1}(t)\rangle$
    $\geq\|g_{1}(t)\|^{2}-\|g_{1}(t)\|\,\|g_{1}(t+\delta t)-g_{1}(t)\|$
    $\geq\|g_{1}(t)\|^{2}-\|g_{1}(t)\|L_{1}|\delta t|.$
    Choose
    (81) $R_{t}^{(1)}:=\frac{\|g_{1}(t)\|}{2L_{1}}>0,$
    Then, whenever $|\delta t|\leq R_{t}^{(1)}$ and $t,t+\delta t\in[0,T]$, the above inequality implies
    (82) $\langle g_{1}(t),g_{1}(t+\delta t)\rangle$ $\geq\|g_{1}(t)\|^{2}-\|g_{1}(t)\|L_{1}R_{t}^{(1)}$
    $=\frac{1}{2}\|g_{1}(t)\|^{2}>0.$
  • If $\|g_{1}(t)\|=0$, then for any $\delta t$ we have
    (83) $\langle g_{1}(t),g_{1}(t+\delta t)\rangle=\langle\mathbf{0},g_{1}(t+\delta t)\rangle=0\geq 0.$
    so the desired inequality holds trivially.
Therefore, for each $t\in[0,T]$, there exists a radius $R_{t}^{(1)}>0$ such that
(84) $\left\langle g_{1}(t),g_{1}(t+\delta t)\right\rangle\geq 0\quad\text{whenever}\quad|\delta t|\leq R_{t}^{(1)},\ \ 0\leq t,\ t+\delta t\leq T.$
Similarly, assume that $\partial^{2}r_{\theta_{2}}/(\partial\theta_{2}\,\partial t)$ is bounded on $[0,T]$, i.e., there exists $L_{2}>0$ such that
(85) $\bigg\|\frac{dg_{2}}{dt}(s)\bigg\|\leq L_{2},\quad\forall s\in[0,T],$
so that $g_{2}$ is $L_{2}$-Lipschitz in $t$ and
(86) $\|g_{2}(t+\delta t)-g_{2}(t)\|\leq L_{2}|\delta t|,\quad\forall t,t+\delta t\in[0,T].$
Fix any $t\in[0,T]$. As before, if $\|g_{2}(t)\|\neq 0$, define
(87) $R_{t}^{(2)}:=\frac{\|g_{2}(t)\|}{2L_{2}}>0,$
and obtain
(88) $\left\langle g_{2}(t),g_{2}(t+\delta t)\right\rangle\geq\frac{1}{2}\|g_{2}(t)\|^{2}>0\quad\text{whenever}\quad|\delta t|\leq R_{t}^{(2)},\ \ 0\leq t,\ t+\delta t\leq T.$
If $\|g_{2}(t)\|=0$, the inner product is again identically zero and the inequality holds trivially:
(89) $\left\langle g_{2}(t),g_{2}(t+\delta t)\right\rangle=\langle\mathbf{0},g_{2}(t+\delta t)\rangle=0\geq 0,$
Hence, for each $t\in[0,T]$, there exists a radius $R_{t}^{(2)}>0$ such that
(90) $\left\langle g_{2}(t),g_{2}(t+\delta t)\right\rangle\geq 0\quad\text{whenever}\quad|\delta t|\leq R_{t}^{(2)},\ \ 0\leq t,\ t+\delta t\leq T.$
Combining the above two parts, for every $t\in[0,T]$ we may set
(91) $R_{t}:=\min\left\{R_{t}^{(1)},\,R_{t}^{(2)}\right\}>0,$
so that
(92) $\left\langle g_{1}(t),g_{1}(t+\delta t)\right\rangle\geq 0\quad\text{and}\quad\left\langle g_{2}(t),g_{2}(t+\delta t)\right\rangle\geq 0,$
hold simultaneously whenever $|\delta t|\leq R_{t}$ and $0\leq t,\ t+\delta t\leq T$.
The only part that is not theoretically guaranteed is the existence of a uniform time region size
(93) $R=\min_{t\in[0,T]}R_{t}>0.$
which would give a single $R$ valid for all time instances, as required in Assumption 3.3.4. In this work, we treat this uniformity as an assumption and support it empirically. The experimental statistics in the following subsection demonstrate that a positive $R$ exists in practice for the considered NeuralMD models. The same reasoning applies to the gated gradient correlation form.
Theorem 3.1 (Gradient correlation in time).
Let Assumption 3.3.4 hold with temporal region size $R>0$. Let $\{\delta t_{i}\}_{i=1}^{k}$ satisfy $|\delta t_{i}|\leq R/3$. Define the temporally averaged quantities
(94) $z^{\mathrm{time}}_{\theta_{1}}(t)=z_{\theta_{1}}(t)+\frac{1}{k}\sum_{i=1}^{k}z_{\theta_{1}}(t+\delta t_{i}),\qquad r^{\mathrm{time}}_{\theta_{2}}(t)=r_{\theta_{2}}(t)+\frac{1}{k}\sum_{i=1}^{k}r_{\theta_{2}}(t+\delta t_{i}),$
Then, for any $\delta t$ satisfying $|\delta t|\leq R/3$, the following inequalities hold:
(95) $G_{z_{\theta_{1}}}(\delta t)\;\leq\;G_{z^{\mathrm{time}}_{\theta_{1}}}(\delta t),\qquad G_{r_{\theta_{2}}}(\delta t)\;\leq\;G_{r^{\mathrm{time}}_{\theta_{2}}}(\delta t).$
Proof 3.2.
We present the argument for $z_{\theta_{1}}$; the case of $r_{\theta_{2}}$ follows identically.
To establish the claim, it suffices to prove that
[…]
Assume $|\delta t|\leq R/3$ and $|\delta t_{i}|\leq R/3$ for all $i=1,\dots,k$. Fix any $t$ such that all involved times lie in $[0,T]$. By the triangle inequality,
(97) $|t-(t+\delta t+\delta t_{j})|$ $=|\delta t+\delta t_{j}|\leq|\delta t|+|\delta t_{j}|\leq\frac{2R}{3}\leq R,$
$|(t+\delta t_{i})-(t+\delta t)|$ $=|\delta t_{i}-\delta t|\leq|\delta t_{i}|+|\delta t|\leq\frac{2R}{3}\leq R,$
$|(t+\delta t_{i})-(t+\delta t+\delta t_{j})|$ $\leq|\delta t_{i}|+|\delta t|+|\delta t_{j}|\leq R.$
Therefore, by Assumption 3.3.4, all inner products corresponding to these pairs are nonnegative
(98) $\left\langle\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t}(t),\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t}(t+\delta t+\delta t_{j})\right\rangle$ $\geq 0,$
$\left\langle\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t}(t+\delta t_{i}),\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t}(t+\delta t)\right\rangle$ $\geq 0,$
$\left\langle\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t}(t+\delta t_{i}),\left.\frac{\partial z_{\theta_{1}}}{\partial\theta_{1}}\right|_{t}(t+\delta t+\delta t_{j})\right\rangle$ $\geq 0$
for all $i,j\in\{1,\dots,k\}$.
Now, consider the following expression
[…]
Thus, we have shown that
$G_{z_{\theta_{1}}}(\delta t)\leq G_{z^{\mathrm{time}}_{\theta_{1}}}(\delta t)\qquad\text{for all }|\delta t|\leq R/3.$
This completes the proof for $z_{\theta_{1}}$. The proof for $r_{\theta_{2}}$ follows in the same way.
Remark 3.3 (Consistency with multiscale decomposition).
For fixed $x\in\Omega$, the functions $G_{z_{\theta_{1}}}(x,\delta t)$ and $G_{r_{\theta_{2}}}(x,\delta t)$ follow the asymptotic forms obtained in the gradient correlation analysis. The leading $\delta t^{2}$ term in both cases is determined by the second spectral moment $\langle k^{2}\rangle_{S,x}$ of the time–frequency density $S_{x}(k,\varepsilon)$. For the amplitude equation, $\langle k^{2}\rangle_{S,x}=O(\varepsilon^{2})$. Hence, the correlation varies only weakly with $\delta t$. For the remainder equation, $\langle k^{2}\rangle_{S,x}=O(\varepsilon^{-2})$. In this case, the correlation decays rapidly. The theorem shows that temporal averaging improves correlation at any fixed time lag $\delta t$. The extent of this improvement is determined by the multiscale frequency structure encoded in $S_{x}(k,\varepsilon)$.
3.4 Training implementation
We have introduced the NeuralMD framework and its algorithmic details. In this section, we implement the training of NeuralMD in two pre-training stages using an unsupervised approach. The training process is summarized in Algorithm 1. Algorithm 1 outlines the training process for NeuralMD.
Algorithm 1 NeuralMD training
0: Sampling the initial data $(\mathbf{x},t)$, learning rate $\eta$, number of epochs NN, steepness parameter $\alpha$, maximum magnitude $\Delta_{max}$, number of perturbation #scale.
1:Pre-training stage I
2:for i=1,,Ni=1,...,N do
3:  Gated gradient flow: $h(t)=(1-\tanh(\alpha(\tilde{t}-\gamma)))/2$
4:  Random time perturbation: $\mathbf{X}_{\text{point}}=\mathcal{P}(\mathbf{x},t),\quad\mathbf{X}_{\text{region}}^{l}=\mathcal{P}(\mathbf{x},t+\delta t_{l}^{\,i})$
5:  Multiscale time mixing: $\mathbf{X}=\mathcal{M}\!\left(\mathbf{X}_{\text{point}},\mathbf{X}_{\text{region}}^{1},\ldots,\mathbf{X}_{\text{region}}^{\text{\#scale}}\right)$
6:  Forward pass through $\mathcal{H}_{\theta^{i}_{1}}:z_{\theta^{i}_{1}}\leftarrow\mathcal{H}_{\theta^{i}_{1}}(\mathbf{X})$
7:  $\mathcal{L}_{\mathrm{Res}}^{z}(\theta^{i}_{1})$, $\mathcal{L}_{\mathrm{Ic}}^{z}(\theta^{i}_{1})$, and $\mathcal{L}_{\mathrm{Bd}}^{z}(\theta^{i}_{1})$ are computed using $z_{\theta^{i}_{1}}$
8:  Compute amplitude loss: $\mathcal{L}^{z}(\theta^{i}_{1})=\lambda_{\mathrm{Res}}\mathcal{L}_{\mathrm{Res}}^{z}(\theta^{i}_{1})+\lambda_{\mathrm{Ic}}\mathcal{L}_{\mathrm{Ic}}^{z}(\theta^{i}_{1})+\lambda_{\mathrm{Bd}}\mathcal{L}_{\mathrm{Bd}}^{z}(\theta^{i}_{1})$
9:  Backpropagate to compute gradients: $\nabla_{\theta^{i}_{1}}{\mathcal{L}}(\theta^{i}_{1})$
10:  Update $\theta^{i}_{1}$ to $\theta^{i+1}_{1}$ using Adam+L-BFGS optimizer
11:  Update the shift parameter $\gamma^{i}$ to $\gamma^{i+1}$ by $\gamma^{i+1}=\gamma^{i}+\eta e^{-\epsilon G_{z_{\theta^{i}_{1}}}}$
12:end for
13:Pre-training stage II
14:for j=1,,Nj=1,...,N do
15:  $h(t),\mathbf{X}$ is the same as the pre-training stage I
16:  Forward pass through $\mathcal{H}_{\theta_{1}^{*}}:z_{\theta^{*}_{1}}\leftarrow\mathcal{H}_{\theta^{*}_{1}}(\mathbf{X})$
17:  Forward pass through $\mathcal{H}_{\theta^{j}_{2}}:r_{\theta^{j}_{2}}\leftarrow\mathcal{H}_{\theta^{j}_{2}}(\mathbf{X})$
18:  $\mathcal{L}_{\mathrm{Res}}^{r}(\theta^{j}_{2})$, $\mathcal{L}_{\mathrm{Ic}}^{r}(\theta^{j}_{2})$, and $\mathcal{L}_{\mathrm{Bd}}^{r}(\theta^{j}_{2})$ are computed using $z_{\theta^{*}_{1}}$ and $r_{\theta^{j}_{2}}$
19:  Compute remainder loss: $\mathcal{L}^{z}(\theta^{j}_{2})=\lambda_{\mathrm{Res}}\mathcal{L}_{\mathrm{Res}}^{z}(\theta^{j}_{2})+\lambda_{\mathrm{Ic}}\mathcal{L}_{\mathrm{Ic}}^{z}(\theta^{j}_{2})+\lambda_{\mathrm{Bd}}\mathcal{L}_{\mathrm{Bd}}^{z}(\theta^{j}_{2})$
20:  Backpropagate to compute gradients: $\nabla_{\theta^{j}_{2}}{\mathcal{L}^{z}}(\theta^{j}_{2})$
21:  Update $\theta^{j}_{2}$ to $\theta^{j+1}_{2}$ using Adam+L-BFGS optimizer
22:  Update the shift parameter $\gamma^{j}$ to $\gamma^{j+1}$ by $\gamma^{j+1}=\gamma^{j}+\eta e^{-\epsilon G_{r_{\theta^{j}_{2}}}}$
23:end for
24:Error criterion
25: Reconstruct the oscillation solution:
$\left\{\begin{aligned} u_{\theta_{1}^{\ast}}(\mathbf{x},t)&=e^{it/\varepsilon^{2}}z_{\theta_{1}^{\ast}}(\mathbf{x},t)+\mathrm{c.c.}\\ u_{(\theta_{1}^{\ast},\theta_{2}^{\ast})}(\mathbf{x},t)&=e^{it/\varepsilon^{2}}z_{\theta_{1}^{\ast}}(\mathbf{x},t)+\mathrm{c.c.}+r_{\theta_{2}^{\ast}}(\mathbf{x},t)\end{aligned}\right$.
26: The prediction solution: $u_{\theta^{\ast}}\leftarrow\mathrm{min}\{\|u_{\theta_{1}^{\ast}}-\hat{u}\|,\|u_{(\theta_{1}^{\ast},\theta_{2}^{\ast})}-\hat{u}\|\}$
27:return Trained models for the amplitude equation and the remainder equation
3.5 Interpretability of NeuralMD
NeuralMD utilizes an MLP-based architecture, grounded in the Universal Approximation Theorem (UAT) [hornik1989mlp], which states that any continuous function $f:\mathbb{R}^{n}\to\mathbb{R}$ can be approximated by a neural network with sufficiently wide hidden layers. For any $\epsilon>0$, there exists a suitable hidden layer width $N(\epsilon)$ such that
(100) $f(x)\approx\sum_{i=1}^{N(\epsilon)}a_{i}\sigma(w_{i}x+b_{i}),$
where $\sigma(\cdot)$ is the activation function (we use tanh, which helps stabilize training and capture the asymptotic behavior of the solution). For deep networks, the structure is expressed as a composition of linear mappings and nonlinear activations, as
(101) $\text{MLP}(x)=(W_{L}\circ\sigma_{L-1}\circ W_{L-1}\circ\cdots\circ\sigma_{1}\circ W_{1})(x).$
However, MLPs concentrate nonlinearity at the nodes, with weight matrices as high-dimensional tensors, leading to highly coupled features that are difficult to interpret in terms of input-output relationships. In NeuralMD, this ”black-box” nature of nonlinear feature mappings complicates the interpretation of the physical meaning of the oscillation-removal process.
To address this, we introduce Kolmogorov–Arnold Networks (KANs) [liu2024kan] as an interpretable alternative. KANs are based on the Kolmogorov–Arnold representation theorem (KAT) [schmidt2021kg], which states that any continuous function $f(x_{1},\ldots,x_{n})$ can be expressed as a finite sum of one-dimensional functions
(102) $f(x_{1},\ldots,x_{n})=\sum_{q=1}^{2n+1}\Phi_{q}\left(\sum_{p=1}^{n}\varphi_{q,p}(x_{p})\right),$
where $\Phi_{q}$ and $\varphi_{q,p}$ are learnable continuous one-dimensional functions. Unlike MLP, which introduces nonlinearity at the nodes, KANs apply nonlinearity at the edges, where each edge corresponds to a learnable one-dimensional function $\varphi_{l,j,i}(\cdot)$, and nodes perform linear summation
(103) $x_{l+1,j}=\sum_{i=1}^{n_{l}}\varphi_{l,j,i}(x_{l,i}),\quad x_{l+1}=\Phi_{l}x_{l}.$
Thus, the entire network can be viewed as a composition of functions:
(104) $\text{KANs}(x)=\left(\Phi_{L-1}\circ\Phi_{L-2}\circ\cdots\circ\Phi_{0}\right)(x),$
allowing each edge to correspond to an explicit one-dimensional function, thereby improving interpretability.
To visualize and symbolize edge functions, KANs use B-splines for parameterization:
(105) $\varphi(x)=w_{b}b(x)+w_{s}\,\text{spline}(x),\quad\text{spline}(x)=\sum_{i}c_{i}B_{i}(x),$
where $b(x)=\text{silu}(x)$, and the control points $c_{i}$ shape the curve of the edge function, explicitly learning the operator terms in the equations. During training, sparse and entropy regularization terms are introduced:
(106) $\mathcal{L}^{z}(\theta_{1})=\lambda_{\mathrm{Res}}\mathcal{L}_{\mathrm{Res}}^{z}(\theta_{1})+\lambda_{\mathrm{Ic}}\mathcal{L}_{\mathrm{Ic}}^{z}(\theta_{1})+\lambda\left(\mu_{1}\sum_{l}\|\Phi_{l}\|_{1}+\mu_{2}\sum_{l}S(\Phi_{l})\right),$
where $S(\Phi_{l})$ is the entropy regularization term, which measures the distribution of information in the edge functions of the $l$-th layer of KANs:
(107) $S(\Phi_{l})=-\sum_{i,j}\frac{\left|\varphi_{l,i,j}\right|}{\sum_{i^{\prime},j^{\prime}}\left|\varphi_{l,i^{\prime},j^{\prime}}\right|}\log\left(\frac{\left|\varphi_{l,i,j}\right|}{\sum_{i^{\prime},j^{\prime}}\left|\varphi_{l,i^{\prime},j^{\prime}}\right|}+\delta\right).$
When each learnable basis function $\Phi_{l,i,j}\in C^{k+1}$, a $k$-th order B-spline approximation exists, and the approximation error is bounded by
(108) $\|f-\left(\Phi_{L-1}^{G}\circ\cdots\circ\Phi_{0}^{G}\right)x\|_{C^{m}}\leq CG^{-(k+1-m)},\quad 0\leq m\leq k,$
where $G$ is the spline density, and $k=3$ for cubic splines.
We integrate KANs into the NeuralMD framework, maintaining the general approximation capability of the operator space while providing structured and symbolic physical interpretability.
4 Numerical experiments
In this section, we demonstrate the numerical performance of the proposed NeuralMD method. We focus on addressing the spectral bias and propagation failure issues in the temporally oscillatory NKGE problem, rather than developing the best possible model. In this section, we first train NeuralMD. Next, we illustrate . Finally, we evaluate the in-distribution generalization and out-of-distribution transfer capabilities of NeuralMD, comparing it with other collocation-based methods. All experiments are conducted on an Nvidia A100-SXM4-80GB GPU. Code and data for the following experiments are available at https://github.com/liangzhangyong/NeuralMD.
4.1 Benchmarks
For all benchmarks, we conduct experiments using three random time regions, denoted as ${\text{\#scale}}=3$, with the corresponding region sizes set to $\{R_{1},R_{2},R_{3}\}=\{1\times 10^{-2},5\times 10^{-2},9\times 10^{-2}\}$, respectively. These regions are designed to represent different temporal scales that are characteristic of the underlying problem. In each region, we perturb the time coordinate by varying the number of perturbations, with values $\{k_{1},k_{2},k_{3}\}=\{3,5,7\}$, ensuring a diverse range of perturbation configurations. The representation feature dimension is set to $d_{\text{model}}=64$, which provides a balanced trade-off between model capacity and computational efficiency. For solving the NKGE in 1D, 2D, and 3D, we follow the experimental protocol outlined in [zhao2023former], training the model using a combination of the Adam optimizer and the L-BFGS optimizer [liu1989limited, kingma2014adam]. The Adam optimizer is initially used for the first 500 iterations to ensure fast convergence in the early training stages, followed by the L-BFGS optimizer for the remaining 500 iterations to refine the solution and achieve more accurate results. The training process spans a total of 1,000 iterations and is implemented in PyTorch [paszke2019torch], a widely used deep learning framework that allows for efficient parallel computation. To assess the model’s performance, we evaluate two key metrics: the Relative L1 Error (rMAE) and the Relative Root Mean Square Error (rRMSE). These metrics provide a robust measure of the model’s accuracy by comparing the predicted solutions against the ground truth. The rMAE quantifies the average absolute error relative to the true solution, while the rRMSE captures the error in terms of the root mean square, giving a clearer sense of the model’s overall predictive accuracy. These metrics are computed over the entire set of evaluation points, allowing for a comprehensive comparison of the model’s performance across different regions.
4.2 Baselines
In addition to vanilla PINNs [raissi2019phy], we also compare NeuralMD with ten PINNs architectures. QRes [bu2021quad], FLS [wong2022learn], CausalPINNs [wang2022resp], PirateNet [wang2024pirat], RoPINNs [wu2024ropinn], ProPINNs [wu2025propinn], PINNsFormer [zhao2023former], SetPINNs [nagda2024set], PINNsMamba [xu2025mamba], MSPINNs [cai2019multi, liu2020multi], and PhasePINNs [cai2020phase] are under the conventional PINNs architecture, where different collocation points are independently optimized. PINNsFormer [zhao2023former], SetPINNs [nagda2024set], and PINNsMamba [xu2025mamba] are based on the Transformer backbone to capture spatiotemporal correlation among PDEs. PirateNet and PINNsFormer are previous state-of-the-art models. In addition, we also integrate sampling strategy R3 [wang2023r3], loss reweighting method [wang2022why], and the latest optimization algorithm RoPINN [wu2024ropinn] to verify that these methods contribute orthogonally to us.
4.3 Model Configuration
In our experiments, we compare NeuralMD with ten baselines. Here are our implementation details for these baselines:
  • For vanilla PINN [raissi2019phy], QRes [bu2021quad], FLS [wong2022learn], and PirateNet [wang2024pirat], we follow the PyTorch implementation of these models provided in the official version. Specifically, vanilla PINN, QRes, and FLS all with 9 layers with 64 hidden channels for the feedforward layer.
  • As for SetPINNs [nagda2024set], PINNsFormer [zhao2023former], and PINNsMamba [xu2025mamba], we implement with only 1 encoder layer, which contains 64 hidden channels for the attention mechanism and 128 hidden channels for the feedforward layer.
  • For RoPINNs [wu2024ropinn] and ProPINNs [wu2025propinn], we use their official code to obtain comparative results. For RoPINNs, we set the initial region size to $10^{-4}$, with 10 past iterations and sample 1 point per region at each iteration for solving the NLGE. Since these points are pre-selected, we do not update their values after sampling and hence do not apply region sampling to them in our experiments. For ProPINNs, we perturb the input dimension $(d+1)$ across 3 scales with sizes 0.03, 0.05, and 0.07. We use uniform sampling with fixed nodes to sample the expanded regions, and the hidden dimension is set to 64.
  • For NeuralMD, we randomly perturb three different time scales to construct multiscale time regions with sizes 0.03, 0.05, and 0.07, and sample 3, 5, and 7 points, respectively. For time regime perturbation sampling, the spatial points remain unchanged, while time points are sampled using a ”tanh” causal gate, with the shift parameter $\gamma$ varying from -0.5 to 1.5. The time-region mixing layer consists of two linear layers with an activation in between, applied only to the time region dimension. This layer first projects the three time scales to 8 and then to 1 after the activation layer. The final projection layer consists of three linear layers with inner activations, and the hidden dimension is set to 64.
4.4 Metrics
The model’s performance is evaluated using two common metrics: the relative L1 error (rMAE) and the relative Root Mean Square Error (rRMSE). For the ground truth $u$ and the predicted solution $u_{\theta}$, these metrics are defined as
(109) $\text{rMAE: }\sqrt{\frac{\sum_{i=1}^{n}\left|u_{\theta}(\mathbf{x}_{i})-u(\mathbf{x}_{i})\right|}{\sum_{i=1}^{n}\left|u(\mathbf{x}_{i})\right|}},\quad\text{rRMSE: }\sqrt{\frac{\sum_{i=1}^{n}\left(u_{\theta}(\mathbf{x}_{i})-u(\mathbf{x}_{i})\right)^{2}}{\sum_{i=1}^{n}\left(u(\mathbf{x}_{i})\right)^{2}}}.$
where $\{\mathbf{x}_{i}\}_{i=1}^{n}$ denotes the set of collocation points used for evaluation.
4.5 1D nonlinear Klein-Gordon equation in whole regime
We evaluate NeuralMD against 12 baselines in the whole regime, conducting experiments in three regions: the relativistic regime ($\varepsilon=0.8$), the transition regime ($\varepsilon=0.5$), and the non-relativistic limit regime ($\varepsilon=0.1,0.01$). We take $d=1$ and $\lambda=1$ in the NKGE (1). Take the initial data as
(110) $\phi_{1}(x)=\frac{e^{-x^{2}}}{\sqrt{\pi}},\quad\phi_{2}(x)=\frac{1}{2}\mathrm{sech}(x^{2})\sin(x),\qquad x\in{\mathbb{R}},$
(111) $\phi_{1}(x)=\frac{3\sin(x)}{e^{x^{2}/2}+e^{-x^{2}/2}},\quad\phi_{2}(x)=\frac{2e^{-x^{2}}}{\sqrt{\pi}},\quad x\in{\mathbb{R}}.,$
As shown in Table 1, for $\varepsilon=0.8$ (low-frequency time oscillation), several models, including PINNs, QRes, PirateNet, MSPINNs, PhasePINNs, and CausalPINNs, fail. In this case, the time oscillation is not the primary cause of failure; instead, the lack of temporal causality is the dominant factor. This suggests that many models struggle with propagation failure in low-frequency time oscillation. MSPINNs and PhasePINNs, which attempt to stretch high-frequency time oscillation into the low-frequency domain, fail to handle time oscillation effectively, even in low-frequency regions. These models do not establish temporal correlation, leading to early propagation failure. The CausalPINNs method, which attempts to establish temporal causality, also fails, indicating its inability to address mild-frequency time oscillation. In contrast, models such as FLS, which use Fourier feature scaling without time correlation, do not experience failure for low-frequency oscillation problems. However, they fail when handling medium-frequency time oscillation at $\varepsilon=0.5$. Several models designed to mitigate propagation failure, including PINNsFormer, SetPINNs, RoPINNs, and PINNsMamba, produce competitive results in the low-frequency region, demonstrating their effectiveness in handling this failure mode. However, they suffer significant accuracy degradation when applied to medium-frequency time oscillation at $\varepsilon=0.5$. The ProPINNs model, with gradient correction across spatiotemporal regions, avoids failure in the medium-frequency regime and maintains competitive accuracy. However, as $\varepsilon$ decreases to 0.1, ProPINNs experiences severe failure, and all models fail with no predictive accuracy.
In contrast, NeuralMD efficiently compensates for the remainder amplitude in the relativistic regime through random time perturbations and multiscale time region mixing, alleviating propagation failure. In the non-relativistic limit regime, NeuralMD effectively removes time oscillation via multiscale decomposition, utilizing a WKB expansion to reconstruct high-frequency time oscillation, thereby overcoming spectral bias. In the transition regime, NeuralMD balances the trade-off between dropping time oscillation and compensating for remainder amplitudes, mitigating both spectral bias and propagation failure. NeuralMD achieves the best performance across the entire regime, with improvements of up to 99.99%, especially in the non-relativistic limit regime ($\varepsilon=0.1,0.01$).
To compare model performance in solving NKGE, we visualize the reference solutions, predicted solutions, and absolute error maps in Figs. 9, 10, and 11. For $\varepsilon=0.8$ (low-frequency time oscillation), most models show competitive predictions (see Fig. 9). However, for $\varepsilon=0.5$ (medium-frequency time oscillation), most models exhibit failure modes (see Fig. 10), and for $\varepsilon=0.1$ (high-frequency time oscillation), all models fail (see Fig. 11). In contrast, NeuralMD outperforms all models, effectively mitigating spectral bias and propagation failure across the entire regime.
Table 1: Performance comparison of different PINNs architectures on the whole regime NKGE. Both rMAE and rRMSE are recorded. Smaller values indicate better performance. For clarity, the best result is in bold and the second best is underlined.
Model $\varepsilon=0.8$ $\varepsilon=0.5$ $\varepsilon=0.1$ $\varepsilon=0.01$
$\mathrm{rMAE}$ $\mathrm{rRMSE}$ $\mathrm{rMAE}$ $\mathrm{rRMSE}$ $\mathrm{rMAE}$ $\mathrm{rRMSE}$ $\mathrm{rMAE}$ $\mathrm{rRMSE}$
PINNs [raissi2019phy] 0.811 0.809 0.883 0.938 1.407 1.540 644.918 811.071
QRes [bu2021quad] 0.761 0.904 0.863 0.945 1.943 1.821 1128.042 1215.761
FLS [wong2022learn] 0.022 0.021 0.891 0.940 2.086 2.422 22.665 31.897
PirateNet [wang2024pirat] 0.692 0.789 0.851 0.923 2.789 2.941 1449.254 1678.351
MSPINNs [cai2019multi] 0.766 0.835 0.854 0.921 1.342 1.458 156.662 214.652
PhasePINNs [cai2020phase] 0.673 0.754 0.784 0.855 1.124 1.246 121.890 189.706
CausalPINNs[wang2022resp] 0.534 0.521 1.382 1.198 25.628 17.085 2479.044 2153.672
PINNsFormer [zhao2023former] 0.022 0.021 0.124 0.146 1.028 1.012 30.362 46.366
SetPINNs [nagda2024set] 0.006 0.008 0.107 0.124 1.026 1.020 30.561 45.286
RoPINNs [wu2024ropinn] 0.022 0.021 1.014 0.953 4.184 5.094 59.345 69.012
ProPINNs [wu2025propinn] 0.004 0.005 0.050 0.053 1.417 1.277 449.188 412.412
PINNsMamba [xu2025mamba] 0.027 0.037 0.235 0.264 1.163 1.042 28.562 35.267
NeuralMD 0.002 0.003 0.008 0.010 0.005 0.007 0.006 0.008
Promotion 50.0% 40.0% 84.0% 81.1% 99.5% 99.3% 99.9% 99.9%
Figure 9: The prediction solution of NeuralMD and baselines for $\varepsilon=0.8,T=5.0$.
Figure 10: The prediction solution of NeuralMD and baselines for $\varepsilon=0.5,T=5.0$.
Figure 11: The prediction solution of NeuralMD and baselines for $\varepsilon=0.1,T=5.0$.
4.6 2D nonlinear Klein-Gordon equation in whole regime
In this section, we explore the wave interactions in 2D. We take $d=2$ and $\lambda=1$ in the NKGE (1) and choose the initial data as
(112) $\begin{split}&\phi_{1}(x,y)=\exp{(-(x+2)^{2}-y^{2})}+\exp{(-(x-2)^{2}-y^{2})},\\ &\phi_{2}(x,y)=\exp{(-x^{2}-y^{2})},\qquad(x,y)\in{\mathbb{R}}^{2}.\end{split}$
The problem is solved numerically on a bounded computational domain $\Omega=(-16,16)\times(-16,16)$ with the periodic boundary condition.
Figure 12: The prediction solution of NeuralMD for $\varepsilon=0.8,T=5.0$.
Figure 13: The prediction solution of NeuralMD for $\varepsilon=0.5,T=5.0$.
Figure 14: The prediction solution of NeuralMD for $\varepsilon=0.1,T=5.0$.
4.7 3D nonlinear Klein-Gordon equation in whole regime
In this section, we explore the wave interactions in 3D. We take $d=3$ and $\lambda=1$ in the NKGE (1) and choose the initial data as
(113) $\begin{split}&\phi_{1}(x,y,z)=2\exp{(-x^{2}-2y^{2}-3z^{2})},\\ &\phi_{2}(x,y,z)=\exp{(-(x+0.5)^{2}-y^{2}-z^{2})},\quad(x,y,z)\in{\mathbb{R}}^{3}.\end{split}$
4.8 Dropping time oscillation and reconstruction
To validate the effectiveness of NeuralMD in dropping time oscillation, we conduct experiments in three regions: the relativistic regime ($\varepsilon=0.8$), the transition regime ($\varepsilon=0.5$), and the non-relativistic limit regime ($\varepsilon=0.1$).
First, as shown in Fig. 15, when $\varepsilon=0.8$, NeuralMD effectively drops time oscillation in the slices at different positions, significantly reducing the oscillation amplitude in low-oscillation regions. At this stage, the impact of time oscillation is minimal, and the process is primarily dominated by the remainder amplitude compensation. A notable observation is that while time oscillation remain at each position, the oscillatory behavior is modulated into a smooth curve. Fig. 16 presents the time oscillation reconstruction using the WKB method with remainder terms, showing good reconstruction results at each position slice with a $L^{2}$ relative error (rRMSE) of $0.064$.
Next, when $\varepsilon=0.5$, Fig. 17 shows the time evolution at different positions, where NeuralMD successfully drops time oscillation, and the oscillatory solutions in the transition region are modulated into smooth, non-oscillatory forms. Fig. 18 shows that after time oscillation reconstruction, NeuralMD’s reconstructed solution closely matches the true solution, effectively alleviating both spectral bias and propagation failure modes, with an $L^{2}$ relative error of $0.099$.
Finally, when $\varepsilon=0.1$, the true solution exhibits high-frequency time oscillation at various positions, as shown in Fig. 19. NeuralMD efficiently drops the high-frequency time oscillation. In the no-oscillation predicted solution, NeuralMD successfully reconstructs the time oscillation (see Fig. 20), reducing the $L^{2}$ relative error to $0.069$ and mitigating the spectral bias caused by high-frequency time oscillation.
Figure 15: Results of dropping time oscillation of NeuralMD for $\varepsilon=0.8,T=5.0$.
Figure 16: Results of reconstructing time oscillation of NeuralMD for $\varepsilon=0.8,T=5.0$.
Figure 17: Results of dropping time oscillation of NeuralMD for $\varepsilon=0.5,T=5.0$.
Figure 18: Results of reconstructing time oscillation of NeuralMD for $\varepsilon=0.5,T=5.0$.
Figure 19: Results of dropping time oscillation of NeuralMD for $\varepsilon=0.1,T=5.0$.
Figure 20: Results of reconstructing time oscillation of NeuralMD for $\varepsilon=0.1,T=5.0$.
4.9 Efficiency comparison
[…]
Figure 21: Efficiency comparison for a temporally oscillatory problem under different $\varepsilon$.
4.9.1 Ablations studies
5 Conclusions
In this work, we proposed NeuralMD, a neural multiscale decomposition framework for solving temporally oscillatory nonlinear Klein–Gordon equations (NKGE) uniformly across $\varepsilon\in(0,1]$. NeuralMD adopts a two-stage pretraining strategy: the first stage learns a modulated nonlinear Schrödinger equation with wave operator (NLSW) with mild-frequency time oscillation, while the second stage learns a small-amplitude remainder equation to mitigate spectral bias induced by time oscillation. The full oscillatory NKGE solution is then reconstructed via a WKB expansion, with an error-based criterion deciding whether the remainder contribution needs to be compensated.
To alleviate propagation failure caused by the modulated NLSW and the oscillatory remainder, we introduced a gated gradient correlation metric. We also propose a gated residual sampling strategy that explicitly couples temporal collocation points along the time axis, thereby strengthening the causal structure of the training dynamics. Furthermore, we extended NeuralMD to an interpretable variant by moving nonlinear activations from nodes to edges, so that nodes perform only summation while edges carry learnable activation functions. In this formulation, the mechanisms of dropping time oscillation and dynamically compensating the remainder amplitude become structurally interpretable at the level of the learned operators. Extensive numerical experiments in different spatial dimensions, with initial data of varying regularity and under long-time prediction settings, show that NeuralMD effectively overcomes both spectral bias and propagation failure compared with existing collocation-based methods, and exhibits robust generalization for temporally oscillatory solutions.
Several directions merit further investigation. First, we plan to enhance NeuralMD to support joint training over the entire parameter regime $\varepsilon\in(0,1]$, where the gradients associated with different regimes may exhibit severe conflicts, making this problem particularly challenging. Second, we aim to extend NeuralMD to semi-classical nonlinear Schrödinger equations with simultaneous temporal and spatial oscillation, where dropping spatio-temporal oscillation is expected to be especially beneficial. Finally, we will explore whether the cascaded structure of the WKB expansion can be exploited to design arbitrarily high-order multiscale time integrator (MTI) schemes, in which parts of the high-order parameter space are constructed end-to-end via operator learning.
Acknowledgments
We would like to thank the referees for their insightful comments that greatly improved the paper. The computing for this work was supported by the High Performance Computing Platform at Eastern Institute of Technology, Ningbo.
Appendix A Generalization analysis for NeuralMD
Here, we discuss the generalization error in expectation, which is independent of the point selection, thereby quantifying the error of NeuralMD optimization more rigorously.
The generalization error in expectation of a model trained on dataset $\mathcal{S}$ is defined as
(114) $\begin{split}\mathcal{E}_{\mathrm{gen}}=\big|\mathbb{E}_{\mathcal{S},\mathcal{A}}\left[{\mathcal{L}\left(u_{\mathcal{A}(\mathcal{S})},\Omega\right)}-\mathcal{L}\left(u_{\mathcal{A}(\mathcal{S})},\mathcal{S}\right)\right]\big|,\end{split}$
where $\mathcal{A}$ denotes the training algorithm and $\mathcal{A}(\mathcal{S})$ represents the optimized model parameters.
{assumption}
The loss function $\mathcal{L}$ is $L$-Lipschitz and $\beta$-smooth with respect to model parameters, which means that $\forall\mathbf{x}\in\Omega$ the following inequalities hold:
(115) $\begin{split}\|\mathcal{L}(u_{\theta_{1}},\mathbf{x})-\mathcal{L}(u_{\theta_{2}},\mathbf{x})\|&\leq L\|\theta_{1}-\theta_{2}\|,\\ \ \|\nabla_{\theta}\mathcal{L}(u_{\theta_{1}},\mathbf{x})-\nabla_{\theta}\mathcal{L}(u_{\theta_{2}},\mathbf{x})\|&\leq\beta\|\theta_{1}-\theta_{2}\|.\end{split}$
Lemma A.1 (Convex case).
Given the stochastic gradient method with an update rule as $G_{\alpha,\mathbf{x}}(\theta)=\theta-\alpha\nabla_{\theta}\mathcal{L}(\theta,\mathbf{x})$ and $\mathcal{L}$ is convex in $\theta$, then for $\alpha\leq\frac{2}{\beta}$, we have $\|G_{\alpha,\mathbf{x}}(\theta_{1})-G_{\alpha,\mathbf{x}}(\theta_{2})\|\leq\|\theta_{1}-\theta_{2}\|$.
Proof A.2.
For clarity, we denote $g=\|\nabla_{\theta}\mathcal{L}(\theta_{1},\mathbf{x})-\nabla_{\theta}\mathcal{L}(\theta_{2},\mathbf{x})\|$. Then we have:
(116) $\begin{split}&\|G_{\alpha,\mathbf{x}}(\theta_{1})-G_{\alpha,\mathbf{x}}(\theta_{2})\|^{2}\\ &=\|\theta_{1}-\theta_{2}-\alpha(\nabla_{\theta}\mathcal{L}(\theta_{1},\mathbf{x})-\nabla_{\theta}\mathcal{L}(\theta_{2},\mathbf{x}))\|^{2}\\ &=\|\theta_{1}-\theta_{2}\|^{2}-2\alpha\left(\nabla_{\theta}\mathcal{L}(\theta_{1},\mathbf{x})-\nabla_{\theta}\mathcal{L}(\theta_{2},\mathbf{x})\right)^{\sf T}(\theta_{1}-\theta_{2})+\alpha^{2}g^{2}\\ &\leq\|\theta_{1}-\theta_{2}\|^{2}-\frac{2\alpha}{\beta}g^{2}+\alpha^{2}g^{2}\\ &\leq\|\theta_{1}-\theta_{2}\|^{2}.\quad\quad\quad\quad\ \ \text{($\alpha\leq\frac{2}{\beta}$)}\end{split}$
Lemma A.3 (Non-convex case).
Given the stochastic gradient method with an update rule as $G_{\alpha,\mathbf{x}}(\theta)=\theta-\alpha\nabla_{\theta}\mathcal{L}(\theta,\mathbf{x})$, then we have $\|G_{\alpha,\mathbf{x}}(\theta_{1})-G_{\alpha,\mathbf{x}}(\theta_{2})\|\leq(1+\alpha\beta)\|\theta_{1}-\theta_{2}\|$.
Proof A.4.
This inequality can be easily obtained from the following:
(117) $\begin{split}&\|G_{\alpha,\mathbf{x}}(\theta_{1})-G_{\alpha,\mathbf{x}}(\theta_{2})\|\\ &=\|\theta_{1}-\theta_{2}-\alpha(\nabla_{\theta}\mathcal{L}(\theta_{1},\mathbf{x})-\nabla_{\theta}\mathcal{L}(\theta_{2},\mathbf{x}))\|\\ &=\|\theta_{1}-\theta_{2}\|+\alpha\|\nabla_{\theta}\mathcal{L}(\theta_{1},\mathbf{x})-\nabla_{\theta}\mathcal{L}(\theta_{2},\mathbf{x})\|\\ &\leq(1+\alpha\beta)\|\theta_{1}-\theta_{2}\|.\end{split}$
Theorem A.5 (Gated multiscale time region optimization).
Suppose that the point optimization loss function $\mathcal{L}$ is $L$-Lipschitz and $\beta$-smooth for $\theta$. Let $h(t)$ be a bounded gate function (e.g., $0\leq h(t)\leq 1$) used in the multi-scale temporal mixing in Eq. (72), and let the corresponding gated multi-scale time loss be denoted by $\mathcal{L}$. If we run stochastic gradient descent with step size $\alpha_{t}$ for $T$ iterations based on $\mathcal{L}$, the generalization error in expectation satisfies:
(1) If $\mathcal{L}$ is convex for $\theta$ and $\alpha_{t}\leq\frac{2}{\beta}$, then
$\mathcal{E}_{\mathrm{gen}}\leq\Bigl(1-\rho\Bigr)\frac{2L^{2}}{|\mathcal{S}|}\sum_{t=1}^{T}\alpha_{t},$
where
$\rho=\sum_{l=1}^{\text{\#scale}}\lambda_{l}\,\bar{h}_{l}\,\frac{|\Omega_{t_{l}}|}{|\Omega|},\qquad\bar{h}_{l}=\frac{1}{|\Omega_{t_{l}}|}\int_{\Omega_{t_{l}}}h(\tau)\,\mathrm{d}\tau,$
and $\lambda_{l}$ is the mixing weight of the $l$-th temporal scale (with $\sum_{l}\lambda_{l}=1$), $\Omega_{t_{l}}$ is the temporal region at scale $l$, and $|\Omega|$ is the measure of the whole temporal domain.
(2) If $\mathcal{L}$ is bounded by a constant $C$ for all $\theta,\mathbf{x}$ and is non-convex for $\theta$ with monotonically non-increasing step sizes $\alpha_{t}\leq\frac{1}{\beta t}$, then
$\mathcal{E}_{\mathrm{gen}}\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}-JL\,\rho^{\,2},$
where $J$ is a finite number that depends on the training property at several beginning iterations.
For clarity, we denote the underlying point-wise loss by $\mathcal{L}_{\text{base}}$ and the gated multiscale time loss in Theorem A.5 by $\mathcal{L}_{\text{gate}}$.
Lemma A.6 (Smoothness of gated multiscale time loss).
If $\mathcal{L}_{\text{base}}$ is bounded for all $\theta,\mathbf{x}$ and is convex, $L$-Lipschitz-$\beta$-smooth with respect to model parameters $\theta$, then $\mathcal{L}_{\text{gate}}$ is also bounded for all $\theta,\mathbf{x}$ and convex, $L$-Lipschitz-$\beta$-smooth for $\theta$.
Proof A.7.
By construction, $\mathcal{L}_{\text{gate}}$ is obtained from $\mathcal{L}_{\text{base}}$ via region-based temporal averaging, bounded gating $h(t)$ with $0\leq h(t)\leq 1$, and a convex combination over scales with weights $\{\lambda_{l}\}_{l=1}^{\#\text{scale}}$, $\sum_{l}\lambda_{l}=1$. Hence boundedness, convexity, $L$-Lipschitz continuity and $\beta$-smoothness are preserved (Lemma A.6), and all conditions of Lemma A.1 and Lemma A.3 hold after replacing $\mathcal{L}$ by $\mathcal{L}_{\text{gate}}$. We denote one SGD step by
$G^{\text{gate}}_{\alpha,\mathbf{x}}(\theta)=\theta-\alpha\nabla_{\theta}\mathcal{L}_{\text{gate}}(\theta,\mathbf{x}).$
We consider two datasets $\mathcal{S}$ and $\mathcal{S}^{\prime}$ differing in exactly one point and denote the corresponding parameter trajectories by $\{\theta_{t}\}_{t=1}^{T}$ and $\{\theta_{t}^{\prime}\}_{t=1}^{T}$. By Lemma LABEL:lemma:sample, it suffices to bound $\mathbb{E}[\|\theta_{T}-\theta_{T}^{\prime}\|]$.
Convex setting
For the convex case, Lemma A.1 implies for $\alpha_{t}\leq 2/\beta$,
(118) $\begin{split}\mathbb{E}\bigl[\|\theta_{t+1}-\theta_{t+1}^{\prime}\|\bigr]&=\Bigl(1-\frac{1}{|\mathcal{S}|}\Bigr)\mathbb{E}\bigl[\|G^{\text{gate}}_{\alpha_{t},\mathbf{x}}(\theta_{t})-G^{\text{gate}}_{\alpha_{t},\mathbf{x}}(\theta_{t}^{\prime})\|\bigr]\\ &\quad+\frac{1}{|\mathcal{S}|}\mathbb{E}\bigl[\|G^{\text{gate}}_{\alpha_{t},\mathbf{x}}(\theta_{t})-G^{\text{gate}}_{\alpha_{t},\mathbf{x}^{\prime}}(\theta_{t}^{\prime})\|\bigr]\\ &\leq\mathbb{E}\bigl[\|\theta_{t}-\theta_{t}^{\prime}\|\bigr]+\frac{1}{|\mathcal{S}|}\mathbb{E}\bigl[\|G^{\text{gate}}_{\alpha_{t},\mathbf{x}}(\theta_{t})-G^{\text{gate}}_{\alpha_{t},\mathbf{x}^{\prime}}(\theta_{t}^{\prime})\|\bigr].\end{split}$
The gated multi-scale temporal gradient at $(\mathbf{x},t)$ can be written as
$\nabla_{\theta}\mathcal{L}_{\text{gate}}(\theta,\mathbf{x},t)=\sum_{l=1}^{\#\text{scale}}\lambda_{l}\frac{1}{|\Omega_{t_{l}}|}\int_{\Omega_{t_{l}}}h(\tau)\,\nabla_{\theta}\mathcal{L}_{\text{base}}(\theta,\mathbf{x},\tau)\,\mathrm{d}\tau.$
For two points $(\mathbf{x},t)$ and $(\mathbf{x}^{\prime},t^{\prime})$, splitting each $\Omega_{t_{l}}$ into the overlapping and non-overlapping parts and using the $L$-Lipschitz property as in the proof of Theorem LABEL:theorem:region_opt, one obtains
(119) $\begin{split}&\mathbb{E}\left[\left\|G^{\mathrm{gate}}_{\alpha_{t},t}(\theta_{t})G^{\mathrm{gate}}_{\alpha_{t},t^{\prime}}(\theta_{t}^{\prime})\right\|\right]\\ &=\mathbb{I}\!\left(1-\rho\right)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\left\|G^{\mathrm{gate}}_{\alpha_{t},t}(\theta_{t})-G^{\mathrm{gate}}_{\alpha_{t},t^{\prime}}(\theta_{t}^{\prime})\right\|\right]+\mathbb{I}\!\left(\rho\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\left\|G^{\mathrm{gate}}_{\alpha_{t},t}(\theta_{t})-G^{\mathrm{gate}}_{\alpha_{t},t^{\prime}}(\theta_{t}^{\prime})\right\|\right]\\[2.84526pt] &\leq\mathbb{I}\!\left(1-\rho\right)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]+\mathbb{I}\!\left(\rho\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\Bigg[\Bigg\|\theta_{t}-\theta_{t}^{\prime}\\ &\quad-\alpha_{t}\sum_{l=1}^{\text{\#scale}}\lambda_{l}\bigg(\frac{1}{\widetilde{|\Omega_{t_{l}}|}}\int_{\Omega_{t_{l}}}\!h(\tau)\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\tau)\,\mathrm{d}\tau-\frac{1}{\widetilde{|\Omega_{t_{l}}|}}\int_{\Omega_{t_{l}}}\!h(\tau)\nabla_{\theta}\mathcal{L}(u_{\theta_{t}^{\prime}},\tau)\,\mathrm{d}\tau\bigg)\Bigg\|\Bigg]\\ &\leq\mathbb{I}\!\left(1-\rho\right)\mathbb{E}_{\Omega_{\mathrm{in}}=0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+2\alpha_{t}L\right]+\mathbb{I}\!\left(\rho\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\|\theta_{t}-\theta_{t}^{\prime}\|+\left(1-\frac{\widetilde{|\Omega_{\mathrm{in}}|}}{\widetilde{|\Omega_{\mathrm{time}}|}}\right)2\alpha_{t}L\right]\\ &\leq\mathbb{E}\left[\|\theta_{t}\theta_{t}^{\prime}\|\right]+2\alpha_{t}L-\mathbb{I}\!\left(\rho\right)\mathbb{E}_{\Omega_{\mathrm{in}}>0}\left[\frac{\widetilde{|\Omega_{\mathrm{in}}|}}{\widetilde{|\Omega_{\mathrm{time}}|}}2\alpha_{t}L\right]\\ &\leq\mathbb{E}\left[\|\theta_{t}-\theta_{t}^{\prime}\|\right]+2\alpha_{t}L\left(1-\rho\right).\end{split}$
where
$\rho=\sum_{l=1}^{\#\text{scale}}\lambda_{l}\,\bar{h}_{l}\,\frac{|\Omega_{t_{l}}|}{|\Omega|},\qquad\bar{h}_{l}=\frac{1}{|\Omega_{t_{l}}|}\int_{\Omega_{t_{l}}}h(\tau)\,\mathrm{d}\tau,$
plays the same role as $\frac{|\Omega_{r}|}{|\Omega|}$ in Theorem LABEL:theorem:region_opt.
Substituting Eq. (118) into Eq. (118) gives
$\mathbb{E}\bigl[\|\theta_{t+1}-\theta_{t+1}^{\prime}\|\bigr]\leq\mathbb{E}\bigl[\|\theta_{t}-\theta_{t}^{\prime}\|\bigr]+\frac{2\alpha_{t}L}{|\mathcal{S}|}(1-\rho).$
Since $\theta_{0}=\theta_{0}^{\prime}$, summing over $t=1,\ldots,T$ yields
$\mathbb{E}\bigl[\|\theta_{T}-\theta_{T}^{\prime}\|\bigr]\leq\frac{2L(1-\rho)}{|\mathcal{S}|}\sum_{t=1}^{T}\alpha_{t}.$
Using the $L$-Lipschitzness of $\mathcal{L}_{\text{gate}}$ and Lemma LABEL:lemma:sample then gives
$\mathcal{E}_{\mathrm{gen}}\leq(1-\rho)\frac{2L^{2}}{|\mathcal{S}|}\sum_{t=1}^{T}\alpha_{t},$
which proves the convex case.
Non-convex setting
For the non-convex case, let $\delta_{t}=\|\theta_{t}-\theta_{t}^{\prime}\|$ and fix $t_{0}\in\{1,\ldots,|\mathcal{S}|\}$. As in Eq. (LABEL:eqn:non_convex_summary),
$\mathbb{E}\bigl[|\mathcal{L}_{\text{gate}}(u_{\theta_{T}},\mathbf{x})-\mathcal{L}_{\text{gate}}(u_{\theta_{T}^{\prime}},\mathbf{x})|\bigr]\leq\frac{Ct_{0}}{|\mathcal{S}|}+L\,\mathbb{E}\bigl[\|\theta_{T}-\theta_{T}^{\prime}\|\mid\delta_{t_{0}}=0\bigr].$
Lemma A.3 applied to $\mathcal{L}_{\text{gate}}$ and the same argument as in the non-convex proof of Theorem A.5 (with $\rho$ in place of $|\Omega_{r}|/|\Omega|$) yield, for $\alpha_{t}\leq 1/(\beta t)$,
(120) $\mathbb{E}\bigl[\delta_{t+1}\mid\delta_{t_{0}}=0\bigr]\leq\Bigl(1+\frac{1}{t}-\frac{1-\rho}{t|\mathcal{S}|}\Bigr)\mathbb{E}[\delta_{t}]+\frac{2L}{\beta t|\mathcal{S}|}(1-\rho).$
Solving the recursion as in Eq. (120), there exists a finite constant $J>0$ (depending only on the first few iterations) such that
$\mathbb{E}\bigl[\|\theta_{T}-\theta_{T}^{\prime}\|\mid\delta_{t_{0}}=0\bigr]\leq\frac{2L}{\beta(|\mathcal{S}|-1)}\Bigl(\frac{T}{t_{0}}-1\Bigr)-J\rho^{2}.$
Setting $t_{0}=1$ and substituting into the previous inequality gives
$\mathbb{E}\bigl[|\mathcal{L}_{\text{gate}}(u_{\theta_{T}},\mathbf{x})-\mathcal{L}_{\text{gate}}(u_{\theta_{T}^{\prime}},\mathbf{x})|\bigr]\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}-JL\rho^{2}.$
Finally, Lemma A.6 implies
$\mathcal{E}_{\mathrm{gen}}\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}-JL\rho^{2},$
which proves the non-convex case of Theorem A.5.
Appendix B Convergence Rate of NeuralMD
Theorem B.1 (Convergence Rate).
Let $h(t)$ be a bounded gate function in the multiscale temporal mixing in (72), and $\mathcal{L}$ the corresponding gated multiscale loss. Assume there exists a constant $H>0$ such that, for all $\bm{v}$ and $\mathbf{x}\in\Omega$,
$\left|\bm{v}^{\sf T}\nabla_{\theta}\mathcal{L}(u_{\theta},\mathbf{x})\bm{v}\right|\leq H\|\bm{v}\|^{2}.$
With step size $\alpha_{t}=1/\sqrt{t+1}$ for $T$ iterations, the stochastic gradient descent method with a Monte Carlo approximation of $\mathcal{L}$ converges at the rate
(121) $\mathbb{E}\left[\left\|\nabla_{\theta}\mathcal{L}(u_{\theta},\mathbf{x})\right\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$
Proof B.2.
By the definition of the multiscale temporal mixing in (72), the gated multiscale time region loss can be written as
(122) $\begin{split}\mathcal{L}(u_{\theta},\bm{x})&=\sum_{l=1}^{\#\text{scale}}\lambda_{l}\,\frac{1}{|\Omega_{t_{l}}|}\int_{\Omega_{t_{l}}}h(\tau)\,\mathcal{L}_{\text{pt}}\bigl(u_{\theta},\bm{x}(\tau)\bigr)\,\mathrm{d}\tau,\end{split}$
where $\mathcal{L}_{\text{pt}}$ denotes the underlying pointwise loss, which is $L$-Lipschitz in $\theta$. Introduce a random variable $\zeta=(l,\tau)$ with distribution
$\mathbb{P}(l=l_{0})=\lambda_{l_{0}},\qquad\tau\sim\text{Unif}(\Omega_{t_{l_{0}}})\ \text{given }l=l_{0}.$
Define the gated stochastic loss
(123) $\ell_{g}(u_{\theta},\bm{x};\zeta):=h(\tau)\,\mathcal{L}_{\text{pt}}\bigl(u_{\theta},\bm{x}(\tau)\bigr),$
so that
(124) $\mathcal{L}(u_{\theta},\bm{x})=\mathbb{E}_{\zeta}\bigl[\ell_{g}(u_{\theta},\bm{x};\zeta)\bigr],\qquad\nabla_{\theta}\mathcal{L}(u_{\theta},\bm{x})=\mathbb{E}_{\zeta}\bigl[\nabla_{\theta}\ell_{g}(u_{\theta},\bm{x};\zeta)\bigr].$
Because $0\leq h(\tau)\leq 1$ and $\mathcal{L}_{\text{pt}}$ is $L$-Lipschitz, we have
(125) $\bigl\|\nabla_{\theta}\ell_{g}(u_{\theta},\bm{x};\zeta)\bigr\|=|h(\tau)|\,\bigl\|\nabla_{\theta}\mathcal{L}_{\text{pt}}(u_{\theta},\bm{x}(\tau))\bigr\|\leq L.$
In the Monte Carlo-based stochastic gradient descent, the update at iteration $t$ is
(126) $\theta_{t+1}=\theta_{t}-\alpha_{t}\nabla_{\theta}\ell_{g}(u_{\theta_{t}},\bm{x};\zeta_{t}),$
where $\zeta_{t}$ is an independent sample from the above distribution. Using a Taylor expansion, there exists $\bm{x}^{\prime}$ such that
(127) $\begin{split}\mathcal{L}(u_{\theta_{t+1}},\bm{x})&=\mathcal{L}\bigl(u_{\theta_{t}}-\alpha_{t}\nabla_{\theta}\ell_{g}(u_{\theta_{t}},\bm{x};\zeta_{t}),\bm{x}\bigr)\\ &=\mathcal{L}(u_{\theta_{t}},\bm{x})-\alpha_{t}\nabla_{\theta}\ell_{g}(u_{\theta_{t}},\bm{x};\zeta_{t})^{\sf T}\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\bm{x})\\ &\quad+\frac{1}{2}\bigl(\alpha_{t}\nabla_{\theta}\ell_{g}(u_{\theta_{t}},\bm{x};\zeta_{t})\bigr)^{\sf T}\nabla_{\theta}^{2}\mathcal{L}(u_{\theta_{t}},\bm{x}^{\prime})\bigl(\alpha_{t}\nabla_{\theta}\ell_{g}(u_{\theta_{t}},\bm{x};\zeta_{t})\bigr).\end{split}$
By the assumption
$\left|\bm{v}^{\sf T}\nabla_{\theta}^{2}\mathcal{L}(u_{\theta},\bm{x})\bm{v}\right|\leq H\|\bm{v}\|^{2},\qquad\forall\,\bm{v},\ \forall\,\bm{x}\in\Omega,$
and using $\bigl\|\nabla_{\theta}\ell_{g}(u_{\theta_{t}},\bm{x};\zeta_{t})\bigr\|\leq L$, we obtain
(128) $\begin{split}\mathcal{L}(u_{\theta_{t+1}},\bm{x})&\leq\mathcal{L}(u_{\theta_{t}},\bm{x})-\alpha_{t}\nabla_{\theta}\ell_{g}(u_{\theta_{t}},\bm{x};\zeta_{t})^{\sf T}\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\bm{x})+\frac{\alpha_{t}^{2}L^{2}H}{2}.\end{split}$
Taking expectations with respect to $\zeta_{t}$ on both sides and using
$\mathbb{E}_{\zeta_{t}}\bigl[\nabla_{\theta}\ell_{g}(u_{\theta_{t}},\bm{x};\zeta_{t})\bigr]=\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\bm{x}),$
we obtain
(129) $\begin{split}\mathbb{E}\bigl[\mathcal{L}(u_{\theta_{t+1}},\bm{x})\bigr]&\leq\mathbb{E}\Bigl[\mathcal{L}(u_{\theta_{t}},\bm{x})-\alpha_{t}\nabla_{\theta}\ell_{g}(u_{\theta_{t}},\bm{x};\zeta_{t})^{\sf T}\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\bm{x})+\frac{\alpha_{t}^{2}L^{2}H}{2}\Bigr]\\ &=\mathbb{E}\bigl[\mathcal{L}(u_{\theta_{t}},\bm{x})\bigr]-\alpha_{t}\mathbb{E}\Bigl[\bigl\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\bm{x})\bigr\|^{2}\Bigr]+\frac{\alpha_{t}^{2}L^{2}H}{2}.\end{split}$
Rearranging terms and summing over $t=0,\dots,T-1$, we have
(130) $\begin{split}\sum_{t=0}^{T-1}\alpha_{t}\mathbb{E}\Bigl[\bigl\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\bm{x})\bigr\|^{2}\Bigr]&\leq\sum_{t=0}^{T-1}\Bigl(\mathbb{E}\bigl[\mathcal{L}(u_{\theta_{t}},\bm{x})\bigr]-\mathbb{E}\bigl[\mathcal{L}(u_{\theta_{t+1}},\bm{x})\bigr]\Bigr)+\sum_{t=0}^{T-1}\frac{\alpha_{t}^{2}L^{2}H}{2}\\ &\leq\mathcal{L}(u_{\theta_{0}},\bm{x})-\mathcal{L}(u_{\theta_{T}},\bm{x})+\frac{L^{2}H}{2}\sum_{t=0}^{T-1}\alpha_{t}^{2}\\ &\leq\mathcal{L}(u_{\theta_{0}},\bm{x})-\mathcal{L}(u_{\ast},\bm{x})+\frac{L^{2}H}{2}\sum_{t=0}^{T-1}\alpha_{t}^{2},\end{split}$
where $u_{\ast}$ denotes a global minimizer of $\mathcal{L}(\cdot,\bm{x})$.
As in the region optimization proof, we now consider a random stopping time $\tau$. For $t=0,\dots,T-1$ set
(131) $\mathbb{P}(\tau=t)=\frac{\alpha_{t}}{\sum_{k=0}^{T-1}\alpha_{k}}.$
Then
(132) $\begin{split}\mathbb{E}\Bigl[\bigl\|\nabla_{\theta}\mathcal{L}(u_{\theta_{\tau}},\bm{x})\bigr\|^{2}\Bigr]&=\Bigl(\sum_{t=0}^{T-1}\alpha_{t}\Bigr)^{-1}\sum_{t=0}^{T-1}\alpha_{t}\mathbb{E}\Bigl[\bigl\|\nabla_{\theta}\mathcal{L}(u_{\theta_{t}},\bm{x})\bigr\|^{2}\Bigr]\\ &\leq\Bigl(\sum_{t=0}^{T-1}\alpha_{t}\Bigr)^{-1}\Bigl(\mathcal{L}(u_{\theta_{0}},\bm{x})-\mathcal{L}(u_{\ast},\bm{x})+\frac{L^{2}H}{2}\sum_{t=0}^{T-1}\alpha_{t}^{2}\Bigr).\end{split}$
With $\alpha_{t}=\frac{1}{\sqrt{t+1}}$, we have
$\sum_{t=0}^{T-1}\alpha_{t}\gtrsim 2\sqrt{T},\qquad\sum_{t=0}^{T-1}\alpha_{t}^{2}=\sum_{t=0}^{T-1}\frac{1}{t+1}\leq\log(T+1),$
and hence
(133) $\begin{split}\mathbb{E}\Bigl[\bigl\|\nabla_{\theta}\mathcal{L}(u_{\theta_{\tau}},\bm{x})\bigr\|^{2}\Bigr]&\lesssim(2\sqrt{T})^{-1}\Bigl(\mathcal{L}(u_{\theta_{0}},\bm{x})-\mathcal{L}(u_{\ast},\bm{x})+\frac{L^{2}H}{2}\log(T+1)\Bigr)\\ &=\mathcal{O}\Bigl(\frac{1}{\sqrt{T}}\Bigr).\end{split}$
This yields the claimed convergence rate. The dependence on the gate function $h(t)$ and the multiscale weights $\{\lambda_{l}\}$ enters only through the definition of $\mathcal{L}$ and the associated constants (e.g., via the effective gating strength $\rho$ in Theorem A.5), while the order $\mathcal{O}(T^{-1/2})$ with respect to $T$ remains unchanged.
Appendix C Gradient estimation error of NeuralMD
Theorem C.1 (Gradient estimation error).
Let the pointwise loss be $\mathcal{L}(u_{\theta},\bm{x},t)$ and define the gated multiscale time-region objective
$\mathcal{L}^{\mathrm{gms}}(u_{\theta},\bm{x},t):=\mathbb{E}_{\delta\sim\pi}\!\Big[h(t+\delta t)\,\mathcal{L}(u_{\theta},\bm{x},t+\delta t)\Big],$
where $\delta t$ is sampled by first choosing $l\sim\mathrm{Cat}(\{\lambda_{l}\})$ and then sampling $\delta t\sim U(\Omega_{t_{l}})$. Its gradient is $\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}=\mathbb{E}_{\delta\sim\pi}[h(t+\delta t)\nabla_{\theta}\mathcal{L}(u_{\theta},\bm{x},t+\delta t)]$. For the one-sample Monte Carlo estimator
$\widehat{g}(\theta;\bm{x},t):=h(t+\delta t)\,\nabla_{\theta}\mathcal{L}(u_{\theta},\bm{x},t+\delta t),$
the gradient estimation error satisfies
$\mathbb{E}_{\delta\sim\pi}\!\Big[\big\|\widehat{g}(\theta;\bm{x},t)-\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}(u_{\theta},\bm{x},t)\big\|^{2}\Big]^{\frac{1}{2}}=\big\|\sigma_{\delta\sim\pi}(\widehat{g}(\theta;\bm{x},t))\big\|.$
Proof C.2.
By the definition of the gated multiscale time-region objective,
$\mathcal{L}^{\mathrm{gms}}(u_{\theta},\bm{x},t)=\mathbb{E}_{\delta t\sim\pi}\!\Big[h(t+\delta t)\,\mathcal{L}(u_{\theta},\bm{x},t+\delta t)\Big],$
where $\delta t$ follows the mixture distribution induced by $l\sim\mathrm{Cat}(\{\lambda_{l}\})$ and $\delta t\sim U(\Omega_{t_{l}})$. Taking gradient w.r.t. $\theta$ and exchanging $\nabla_{\theta}$ with $\mathbb{E}$ yields
$\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}(u_{\theta},\bm{x},t)=\mathbb{E}_{\delta t\sim\pi}\!\Big[h(t+\delta t)\,\nabla_{\theta}\mathcal{L}(u_{\theta},\bm{x},t+\delta t)\Big],$
Recall the one-sample Monte Carlo estimator
$\widehat{g}(\theta;\bm{x},t)=h(t+\delta t)\,\nabla_{\theta}\mathcal{L}(u_{\theta},\bm{x},t+\delta t),\qquad\delta t\sim\pi,$
Then the estimation error satisfies
$\mathbb{E}_{\delta t\sim\pi}\!\Big[\big\|\widehat{g}(\theta;\bm{x},t)-\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}(u_{\theta},\bm{x},t)\big\|^{2}\Big]^{\frac{1}{2}}$
$=\mathbb{E}_{\delta t\sim\pi}\!\Big[\big\|h(t+\delta t)\nabla_{\theta}\mathcal{L}(u_{\theta},\bm{x},t+\delta t)-\mathbb{E}_{\delta t\sim\pi}[h(t+\delta t)\nabla_{\theta}\mathcal{L}(u_{\theta},\bm{x},t+\delta t)]\big\|^{2}\Big]^{\frac{1}{2}}$
$=\big\|\sigma_{\delta t\sim\pi}\!\big(\widehat{g}(\theta;\bm{x},t)\big)\big\|.$
where $\sigma_{\delta t\sim\pi}(\cdot)$ denotes the standard deviation of the random gradient under $\delta t\sim\pi$. This proves the theorem.
Theorem C.3 (NeuralMD with gradient estimation error).
Assume the pointwise loss $\mathcal{L}(u_{\theta},\bm{x},t)$ is $L$-Lipschitz and $\beta$-smooth w.r.t. $\theta$. Let $\delta t\sim\pi$ be the multiscale mixture: choose $l\sim\mathrm{Cat}(\{\lambda_{l}\})$ and then sample $\delta t\sim U(\Omega_{t_{l}})$. Define the gated multiscale time region objective
$\mathcal{L}^{\mathrm{gms}}(u_{\theta},\bm{x},t):=\mathbb{E}_{\delta t\sim\pi}\!\Big[h(t+\delta t)\,\mathcal{L}(u_{\theta},\bm{x},t+\delta t)\Big],\quad g:=\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}.$
At iteration $k$, use the one-sample gradient estimator
$\widehat{g}_{k}:=h(t+\delta t_{k})\,\nabla_{\theta}\mathcal{L}(u_{\theta},\bm{x},t+\delta t_{k}),\qquad\delta t_{k}\sim\pi,$
and assume a uniform bound on the gradient estimation error
$\mathcal{E}_{\mathrm{gms,grad}}:=\max_{k\leq T}\|\widehat{g}_{k}-g\|.$
Let
$\rho:=\sum_{l=1}^{\#\mathrm{scale}}\lambda_{l}\,\bar{h}_{l}\,\frac{|\Omega_{t_{l}}|}{|\Omega|},\qquad\bar{h}_{l}:=\frac{1}{|\Omega_{t_{l}}|}\int_{\Omega_{t_{l}}}h(\tau)\,d\tau.$
Run SGD for $T$ steps with step sizes $\{\alpha_{k}\}_{k=1}^{T}$ based on $\widehat{g}_{k}$. Then:
(1) If $\mathcal{L}$ is convex in $\theta$ and $\alpha_{k}\leq 2/\beta$,
$\mathcal{E}_{\mathrm{gen}}\leq\Big((1-\rho)L+\mathcal{E}_{\mathrm{gms,grad}}\Big)\frac{2L}{|\mathcal{S}|}\sum_{k=1}^{T}\alpha_{k}.$
(2) If $\mathcal{L}$ is bounded by $C$, non-convex in $\theta$, and $\alpha_{k}$ is non-increasing with $\alpha_{k}\leq 1/(\beta k)$,
$\mathcal{E}_{\mathrm{gen}}\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}-J^{\prime}L\rho^{2}+J^{\prime}\mathcal{E}_{\mathrm{gms,grad}}(1+\rho),$
where $J^{\prime}$ is a finite constant depending on the first few iterations.
Convex setting
Proof C.4.
Let $\mathcal{S}$ and $\mathcal{S}^{\prime}$ be two training sets differing by one sample. Let $\{\theta_{k}\}$ and $\{\theta^{\prime}_{k}\}$ be the SGD iterates trained on $\mathcal{S}$ and $\mathcal{S}^{\prime}$, respectively. Define the one-step update map using the one-sample gated multiscale gradient
$G^{\mathrm{approx}}_{\alpha_{k},z}(\theta)$ $:=\theta-\alpha_{k}\,\nabla_{\theta}\mathcal{L}^{\mathrm{gms,approx}}(\theta;z),\quad$
$\nabla_{\theta}\mathcal{L}^{\mathrm{gms,approx}}(\theta;z)$ $=h(t+\delta t)\,\nabla_{\theta}\mathcal{L}(\theta;z,t+\delta t),\ \ \delta t\sim\pi,$
where $z=(\bm{x},t)$. Similarly denote the true gated multiscale gradient update
$G^{\mathrm{gms}}_{\alpha_{k},z}(\theta)$ $:=\theta-\alpha_{k}\,\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}(\theta;z),\quad$
$\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}(\theta;z)$ $=\mathbb{E}_{\delta t\sim\pi}\!\left[h(t+\delta t)\nabla_{\theta}\mathcal{L}(\theta;z,t+\delta t)\right].$
At step $k$, by conditioning on whether the sampled data point is the differing one
$\begin{split}\mathbb{E}\!\left[\|\theta_{k+1}-\theta^{\prime}_{k+1}\|\right]&=\Bigl(1-\frac{1}{|\mathcal{S}|}\Bigr)\mathbb{E}\!\left[\|G^{\mathrm{approx}}_{\alpha_{k},z}(\theta_{k})-G^{\mathrm{approx}}_{\alpha_{k},z}(\theta^{\prime}_{k})\|\right]\\ &\qquad+\frac{1}{|\mathcal{S}|}\mathbb{E}\!\left[\|G^{\mathrm{approx}}_{\alpha_{k},z}(\theta_{k})-G^{\mathrm{approx}}_{\alpha_{k},z^{\prime}}(\theta^{\prime}_{k})\|\right]\\ &\leq\Bigl(1-\frac{1}{|\mathcal{S}|}\Bigr)\mathbb{E}\!\left[\|\theta_{k}-\theta^{\prime}_{k}\|\right]+\frac{1}{|\mathcal{S}|}\mathbb{E}\!\left[\|G^{\mathrm{approx}}_{\alpha_{k},z}(\theta_{k})-G^{\mathrm{approx}}_{\alpha_{k},z^{\prime}}(\theta^{\prime}_{k})\|\right].\end{split}$
We bound the second term by adding and subtracting the true gated multiscale gradients
$\begin{split}&\mathbb{E}\!\left[\|G^{\mathrm{approx}}_{\alpha_{k},z}(\theta_{k})-G^{\mathrm{approx}}_{\alpha_{k},z^{\prime}}(\theta^{\prime}_{k})\|\right]\\ &\leq\mathbb{E}\!\left[\|G^{\mathrm{gms}}_{\alpha_{k},z}(\theta_{k})-G^{\mathrm{gms}}_{\alpha_{k},z^{\prime}}(\theta^{\prime}_{k})\|\right]\\ &\qquad+\alpha_{k}\mathbb{E}\!\left[\|\nabla_{\theta}\mathcal{L}^{\mathrm{gms,approx}}(\theta_{k};z)-\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}(\theta_{k};z)\|\right]\\ &\qquad+\alpha_{k}\mathbb{E}\!\left[\|\nabla_{\theta}\mathcal{L}^{\mathrm{gms,approx}}(\theta^{\prime}_{k};z^{\prime})-\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}(\theta^{\prime}_{k};z^{\prime})\|\right].\end{split}$
For the first term, the convex-case stability argument for gated multiscale time-region optimization (Theorem A.5) implies
$\mathbb{E}\!\left[\|G^{\mathrm{gms}}_{\alpha_{k},z}(\theta_{k})-G^{\mathrm{gms}}_{\alpha_{k},z^{\prime}}(\theta^{\prime}_{k})\|\right]\leq\mathbb{E}\!\left[\|\theta_{k}-\theta^{\prime}_{k}\|\right]+2\alpha_{k}L(1-\rho),$
where $\rho=\sum_{l}\lambda_{l}\bar{h}_{l}\frac{|\Omega_{t_{l}}|}{|\Omega|}$. For the remaining two terms, by the definition of the uniform gradient estimation bound $\mathcal{E}_{\mathrm{gms,grad}}=\max_{k\leq T}\|\widehat{g}_{k}-g\|$, we have
$\mathbb{E}\!\left[\|\nabla_{\theta}\mathcal{L}^{\mathrm{gms,approx}}-\nabla_{\theta}\mathcal{L}^{\mathrm{gms}}\|\right]\leq\mathcal{E}_{\mathrm{gms,grad}},$
Combining the above yields
$\mathbb{E}\!\left[\|G^{\mathrm{approx}}_{\alpha_{k},z}(\theta_{k})-G^{\mathrm{approx}}_{\alpha_{k},z^{\prime}}(\theta^{\prime}_{k})\|\right]\leq\mathbb{E}\!\left[\|\theta_{k}-\theta^{\prime}_{k}\|\right]+2\alpha_{k}\Big(L(1-\rho)+\mathcal{E}_{\mathrm{gms,grad}}\Big).$
Plugging back and unrolling the recursion gives the uniform stability
$\sup_{z}\mathbb{E}\!\left[|\mathcal{L}(u_{\theta_{T}},z)-\mathcal{L}(u_{\theta^{\prime}_{T}},z)|\right]\leq\Big(L(1-\rho)+\mathcal{E}_{\mathrm{gms,grad}}\Big)\frac{2L}{|\mathcal{S}|}\sum_{k=1}^{T}\alpha_{k}.$
which proves the convex-case bound.
Non-convex setting
Proof C.5.
As in the region-opt proof, $\mathcal{L}^{\mathrm{gms,approx}}$ remains $L$-Lipschitz and $\beta$-smooth w.r.t. $\theta$ (since $0\leq h(\cdot)\leq 1$ and the sampling does not change smoothness constants). Let $\delta_{k}:=\mathbb{E}\!\left[\|\theta_{k}-\theta^{\prime}_{k}\|\right]$ and denote $\rho$ as above.
For step sizes $\alpha_{k}\leq\frac{1}{\beta k}$ and conditioning on $\delta_{k_{0}}=0$, we obtain
(134) $\begin{split}\mathbb{E}\!\left[\|\theta_{k+1}-\theta^{\prime}_{k+1}\|\mid\delta_{k_{0}}=0\right]&\leq\Bigl(1-\frac{1}{|\mathcal{S}|}\Bigr)\Bigl(1+\frac{1}{k}\Bigr)\mathbb{E}[\delta_{k}]\\ &\qquad+\frac{1}{|\mathcal{S}|}\mathbb{E}\!\left[\|G^{\mathrm{approx}}_{\alpha_{k},z}(\theta_{k})-G^{\mathrm{approx}}_{\alpha_{k},z^{\prime}}(\theta^{\prime}_{k})\|\right]\\ &\leq\Bigl(1+\frac{1}{k}-\frac{1-\rho}{k|\mathcal{S}|}\Bigr)\mathbb{E}[\delta_{k}]+\frac{2\alpha_{k}}{|\mathcal{S}|}\Big(L(1-\rho)+\mathcal{E}_{\mathrm{gms,grad}}\Big)\\ &\leq\exp\!\Bigl(\frac{1}{k}-\frac{1-\rho}{k|\mathcal{S}|}\Bigr)\mathbb{E}[\delta_{k}]+\frac{2}{\beta k|\mathcal{S}|}\Big(L(1-\rho)+\mathcal{E}_{\mathrm{gms,grad}}\Big).\end{split}$
Similarly to the region-opt proof, when the early-stage condition analogous to $\mathbb{E}(\delta_{k})\leq\frac{2L}{\beta}-\frac{2}{\beta M}\mathcal{E}_{r,\mathrm{grad}}$ holds (with $M$ replaced by $\rho$), we can accumulate Eq. (134) over the first $K^{\prime}$ steps and obtain
$\Delta\leq\sum_{k=k_{0}+1}^{k_{0}+K^{\prime}}\exp\!\Bigl((1-\tfrac{1}{|\mathcal{S}|})\log\tfrac{T}{k}\Bigr)\frac{2L}{\beta k|\mathcal{S}|}\;-\;J^{\prime}L\rho^{2}\;+\;J^{\prime}\mathcal{E}_{\mathrm{gms,grad}}(1+\rho),$
where $J^{\prime}$ is a finite constant depending on the first few iterations (i.e., $k_{0},K^{\prime}$). Proceeding as in Lamma A.1, summing over all $T$ steps yields
$\mathbb{E}\!\left[\|\theta_{T}-\theta^{\prime}_{T}\|\mid\delta_{k_{0}}=0\right]\leq\frac{2L}{\beta(|\mathcal{S}|-1)}\sum_{k=1}^{T-1}\frac{1}{|\mathcal{S}|}\;-\;J^{\prime}\rho^{2}\;+\;J^{\prime}\mathcal{E}_{\mathrm{gms,grad}}(1+\rho).$
Finally, using the standard stability-to-generalization argument, we obtain
$\mathcal{E}_{\mathrm{gen}}\leq\frac{C}{|\mathcal{S}|}+\frac{2L^{2}(T-1)}{\beta(|\mathcal{S}|-1)}-J^{\prime}L\rho^{2}+J^{\prime}\mathcal{E}_{\mathrm{gms,grad}}(1+\rho).$
which proves the non-convex-case bound.
Appendix D Additional results
In this section, we present additional results, including further hyperparameter analysis, new experiments, and additional examples, to supplement the main text.
Appendix E Inverse problem experiments
As we stated in the implementations, NeuralMD can also be applied to tasks with data loss for the inverse problem.
Appendix F Impact of optimizers
In this work, we adhere to established benchmarks and use Adam+L-BFGS for standard evaluations. This section examines the influence of different optimizers on the performance of NeuralMD.
Appendix G Hyperparameter Analysis
This section evaluates the model’s performance across various hyperparameter configurations, including the number of random time perturbations at each scale $(k_{1},k_{2},k_{3})$, the size of the perturbation region $(R_{1},R_{2},R_{3})$, and the number of scales (#scales).
  • Increasing the number of perturbations improves the model’s performance.
  • The perturbation region size is determined by the NKGE property.
  • Incorporating additional scales can enhance the model’s performance.